### v1

# Cut the Deadwood Out: Post-Training Model Purification with Selective Module Substitution

Abstract
The success of DNNs often depends on training with large-scale datasets, but building such datasets is both expensive and challenging. Consequently, public datasets from open-source platforms like HuggingFace have become popular, posing significant risks of data poisoning attacks. Existing backdoor defenses in NLP primarily focus on identifying and removing poisoned samples; however, purifying a backdoored model with these sample-cleaning approaches typically requires expensive retraining. Therefore, we propose Greedy Module-Substitution (GMS), which identifies and substitutes “deadwood” modules (i.e., components critical to backdoor pathways) in a backdoored model to purify it. Our method relaxes the common dependency of prior model purification methods on clean datasets or clean auxiliary models. When applied to RoBERTa-large under backdoor attacks, GMS demonstrates strong effectiveness across various settings, particularly against widely recognized challenging attacks like LWS, achieving a post-purification attack success rate (ASR) of 9.7% on SST-2 compared to 58.8% for the best baseline approach.
Cut the Deadwood Out: Post-Training Model Purification
with Selective Module Substitution
Yao Tong1thanks: Equal contribution., Weijun Li2∗, Xuanli He3, Haolan Zhan4, Qiongkai Xu2thanks: The corresponding author. 1 National University of Singapore, 2 Macquarie University 3 University College London, 4 Monash University tongyao@u.nus.edu, weijun.li1@hdr.mq.edu.au xuanli.he@ucl.ac.uk, haolan.zhan@monash.edu, qiongkai.xu@mq.edu.au
1 Introduction
Deep neural networks (DNNs) have achieved remarkable success in various domains Krizhevsky et al. (2012); Bowman et al. (2016), largely driven by training on large-scale datasets. Given the high cost of collecting and labeling data from scratch, incorporating public datasets from open-source platforms like HuggingFace in training has become a common practice, particularly in NLP Howard and Ruder (2018); Ding et al. (2023). However, this growing reliance on open-source datasets from unaudited platforms raises significant security concerns Yang et al. (2021a); Zhang et al. (2021), especially for commercial use. Malicious actors can exploit this ecosystem by publishing poisoned datasets Xu et al. (2021) to embed backdoors into models for potential benefit. These backdoors allow attackers to manipulate models, causing targeted misbehavior when triggered by specific inputs while operating normally on clean inputs. As a result, they often remain stealthy until their triggers are activated.
Figure 1: Parameter-level editing (left) prunes parameters "most relevant" to backdoor features but leaves many backdoor-related parameters intact within each module, allowing backdoor pathways to potentially persist. In contrast, our module-level substitution (right) effectively disrupts these pathways by completely replacing backdoor-connected modules with counterparts from a proxy model (not required to be clean).
Existing defense strategies against backdoor attacks in NLP primarily focus on backdoor samples detection, such as detecting and removing poisoned samples leveraging their distinct characteristics compared to clean ones Qi et al. (2021a); Yang et al. (2021b); Tran et al. (2018); He et al. (2023), or reverse-engineer backdoor triggers to neutralize their impact Wang et al. (2022, 2023); Xu et al. (2024). However, their effectiveness is often suboptimal Min et al. (2024), especially as backdoor triggers evolve to be stealthier and harder to detect or reverse Qi et al. (2021b). Moreover, while these methods follow a fine-tuning or adversarial training for unlearning Wei et al. (2023) to sanitize poisoned models, they cannot effectively purify a backdoored model without incurring a significant cost of retraining Wu et al. (2022). Then a practical question arises: How can we cost-efficiently eliminate the backdoor features from an existing model after users reported some compromise?
While only a few model purification defenses in NLP exist, they predominantly focus on parameter-level editing. For instances, Fine-mixing Zhang et al. (2022) and Fine-purification Zhang et al. (2023) mix a clean pre-trained language model (PLM) with a backdoored model and rely on fine-tuning with a clean dataset for purification. Similarly, Zhao et al. (2024) prunes attention heads potentially affected by poisoned data and fine-tunes the pruned model using a clean dataset. However, recent studies have observed that backdoor features are encoded as complex pathways in victim models, with significant overlap between backdoor and clean neurons at the module level Jebreel et al. (2023); Li et al. (2024). As illustrated in Figure 1 (Left), parameter-level editing will leave hidden vulnerabilities in the non-pruned connections, thus making the defense effectiveness heavily relies on fine-tuning with a clean dataset to restore clean features Liu et al. (2018), an assumption that is both strong and rarely feasible in practice.
Here, we develop a method that effectively disrupts the model’s backdoor pathways without relying on the assumption of access to a clean dataset. Note that the abundance of public platforms for open-source models (e.g., HuggingFace) offers semantically similar yet independent models for any task domain Semwal et al. (2018). Therefore, we propose a simple yet effective module-level purification defense, Greedy Module-Substitution (GMS), which identifies critical modules for backdoor propagation and purifies them by replacing the parameters with those of a proxy model. While this approach eliminates the need for a clean dataset to restore semantic knowledge, the proxy model is not required to be clean either, since the pathways of two different backdoored models are less likely to be connected (empirically verified in Table 5). The primary contributions of this work are as follows.
  1. 1.
    We propose the first efficient (quadratic time complexity) module-level backdoor purification defense (GMS), which relaxes the strong reliance on clean auxiliary models and clean datasets required by prior defenses, as shown in Table 5 and Table 6.
  2. 2.
    Our method significantly outperforms baseline approaches against challenging backdoor attacks like LWS and HiddenKiller, e.g., on SST-2 victim model, GMS reduces the ASR on LWS to 9.7%, compared to 12.8% for the benign model and 58.8% for the best baseline.
  3. 3.
    Our GMS purification strategy exhibits strong transferability across different attacks. Even the strategy with few module substitutions can effectively defend against other backdoor attacks, making it suitable when the victim dataset is inaccessible (Section 5.3 and Table 7).
2 Related works
Backdoor attack
Backdoor attacks aim to manipulate models to behave normally on benign inputs while exhibiting attacker-controlled behavior when specific triggers are present. They can be categorized into two series Wu et al. (2022): (1) Data-poisoning attacks, exemplified by BadNets Gu et al. (2017), involve tampering with a subset of training data by adding triggers and altering labels, making them practical in real-world scenarios with minimal attacker assumptions Turner et al. (2019); Carlini and Terzis (2022); Goldblum et al. (2023); Li et al. (2021c). Initially explored in image classification, backdoor attacks have since raised significant security concerns in NLP Dai et al. (2019); Qi et al. (2021d, c) with triggers range from misspelled words Chen et al. (2021), rare words Yang et al. (2021a); Kurita et al. (2020) and syntactic structures Qi et al. (2021c) to text styles Pan et al. (2022), posing serious challenges for detection. (2) Training-control attacks, on the other hand, assume complete control over the training process and data Nguyen and Tran (2021); Kurita et al. (2020), with advances such as layer-wise weight poisoning ensuring backdoor persistence even after fine-tuning Li et al. (2021a). In this work, we focus on data-poisoning backdoor attacks in NLP due to their widespread applicability and practical implications.
Backdoor defense
Backdoor defense approaches can be categorized according to their strategies: (1) Backdoor sample detection focuses on identifying poisoned samples in the dataset. The defense approaches usually include detecting and removing backdoor samples, leveraging the assumption that poisoned samples exhibit distinct characteristics compared to clean (benign) ones Qi et al. (2021a); Yang et al. (2021b); Tran et al. (2018); He et al. (2023), or reverse-engineer backdoor triggers to neutralize their impact Wang et al. (2019, 2022, 2023); Xu et al. (2024). These strategies are typically applied during the training phase or involve fine-tuning after removing backdoor samples. (2) Backdoor model purification: These methods aim to eliminate backdoor features from a well-trained model. Common approaches involve merging parameters with other proxy models Arora et al. (2024); Chen et al. (2024a), or fine-tuning or pruning using a limited amount of clean data Wu and Wang (2021); Min et al. (2023); Liu et al. (2018); Zhao et al. (2024); Zhu et al. (2023). These techniques not only reduce training costs but also achieve state-of-the-art defense performance Zhu et al. (2023); Zhao et al. (2024); Arora et al. (2024). Our work advances the second line of research by minimizing the number of auxiliary models to one and reducing dependence on the clean dataset.
3 Preliminary
Modern neural network (NN) architectures can be viewed as layer-wise compositions of functional blocks, where each block serves as a reusable unit performing a well-defined computational operation. Hereby, we define two aspects to systematically describe generic NN architectures:
  1. 1.
    Layer Set ($\mathcal{L}$) is the set of block indices, capturing the architecture’s depth and structure.
  2. 2.
    Module Set ($\mathcal{M}$) is the set of representative functional modules within each block.
Transformer blocks
In this paper, we use Transformer-based architecture Vaswani et al. (2017) as the testbed, given its widespread attention and implementation in NLP.
For each Transformer block, the input will be processed through Attention and Feed-Forward Network (FFN) modules. Attention modules contains four components: $W_{Q}$, $W_{K}$, $W_{V}$, and $W_{O}$, projecting the input into query, key, value, and the final output representation, respectively. The following FFN contains two components: $W_{F}$ for the forward layer and $W_{P}$ for the projection layer, connected by a non-linear activation function.
In total, each Transformer block contains six key functional modules with parameters $W_{Q},W_{K},W_{V},W_{O},W_{F},$ and $W_{P}$. For simplicity and consistency, we adopt notations $Q,K,V,O,F,$ and $P$ to refer to these weight matrices throughout the paper. Accordingly, we define the module set $\mathcal{M}$ for Transformer models as $\{Q,K,V,O,F,P\}$ and layer set =$\mathcal{L}=\{1,2,\ldots,L\}$.
Evaluation metrics
Following previous literature (Qi et al., 2020), we adopt two metrics: (1) Clean Accuracy (CACC) measures the prediction accuracy on clean samples, and (2) Attack Success Rate (ASR) measures the attack accuracy on poisoned test set. A higher CACC reflects better utility, while lower ASR indicates stronger defense.
Figure 2: The pipeline of our method. Step 1: Extract two small proxy datasets from the suspected dataset $\mathcal{D}$ used for computing score in Equation 1. Step 2: Iteratively update a substitution matrix $S$ to greedily maximize the trade-off score between backdoor removal and utility preservation. Step 3: Return purified model $M_{pure}$ corresponding to the best substitution matrix $S_{best}$. For more details, refer to Algorithm 1.
4 Methodology
In Section 4.1, we introduce the problem setting, and in Section 4.2 we present our methods.
4.1 Problem setting
Threat Model
We consider the scenario of data poisoning-based backdoor attacks, where an attacker releases a poisoned dataset $\mathcal{D}$ to embed backdoors into any model trained on this dataset. Starting with a clean dataset $\mathcal{D}_{\text{cl}}=\{(x_{i},y_{i})\}_{i=1}^{|\mathcal{D}|}\subseteq\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ and $\mathcal{Y}$ denote the input and label spaces, respectively, the attacker selects a subset $\mathcal{D}_{\text{sub}}\subset\mathcal{D}_{\text{cl}}$, and then modified each sample $(x,y)\in\mathcal{D}_{\text{sub}}$ into a poisoned sample $(\tilde{x},y_{t})$ where $\tilde{x}=g(x,\Delta)$, the pre-defined trigger is $\Delta$, $g$ is a generation function, and $y_{t}$ is the target label. The poisoned subset is $\mathcal{D}_{\text{poi}}=\{(\tilde{x},y_{t})\mid(x,y)\in\mathcal{D}_{\text{sub}}\}$. Finally, the attackers publish the poisoned dataset constructed as $\mathcal{D}=(\mathcal{D}_{\text{cl}}\setminus\mathcal{D}_{\text{sub}})\cup\mathcal{D}_{\text{poi}}$.
Defense Scenario
The defender (e.g., a model vendor) has already trained a (potentially backdoored) victim model $M_{v}$ on an unauthenticated dataset $\mathcal{D}$. While the defender has access to $\mathcal{D}$, they are unaware of all possible backdoor patterns. [^1] As shown in Section 4.2.2, the full training dataset is not strictly required; a small subset of poisoned data reported by users or automatically detected by algorithms is sufficient for applying our defense. The defender has white-box access to a trained victim model $M_{v}$ and a proxy model $M_{p}$ trained on similar downstream tasks.
4.2 Our method
We aim to disrupt the backdoor connection in victim model $M_{v}$ while preserving the knowledge required for the original task. Models trained on similar tasks tend to capture similar domain-specific knowledge Semwal et al. (2018). Hence, we replace backdoor-connected modules in the $M_{v}$ with counterparts from a proxy model $M_{p}$ that shares similar domain knowledge. Notably, $M_{p}$ does NOT need to be benign, as even a malicious proxy model lacking identical backdoor pathways will fail to restore the disrupted backdoor connection.
Our main algorithm, Greedy Module-Substitution (GMS), is detailed in Section 4.2.3. Section 4.2.1 defines the objective optimized by GMS, and Section 4.2.2 describes the method to construct the proxy datasets used in evaluation.
4.2.1 Objective
The defender’s objective is to identify a purified model $M_{pure}$ from all possible substituted models $M$ that effectively eliminates the backdoor while preserving task utility. This can be formulated as
$\operatorname*{argmax}_{M}\,(1-\alpha)\cdot\delta_{\text{asr}}(M)+\alpha\cdot(1-\delta_{\text{acc}}(M))$ (1)
where $\delta_{\text{asr}}(M)$ measures the extent of backdoor removal by quantifying the reduction in ASR and $\delta_{\text{acc}}(M)$ assesses the impact on task utility by measuring the decrease in CACC. However, since the defender lacks access to the clean dataset and poisoned test set, we adopt the method described in Section 4.2.2 to extract two proxy datasets, $\mathcal{D}_{\text{poison}}$ and $\mathcal{D}_{\text{clean}}$, from the suspected dataset $\mathcal{D}$ for ASR and CACC estimation.
The goal is to maximize $\delta_{\text{asr}}(M)$ while minimizing $\delta_{\text{acc}}(M)$. The hyperparameter $\alpha\in[0,1]$ controls the trade-off between backdoor defense and task utility. A larger $\alpha$ favors utility, while a smaller $\alpha$ prioritizes backdoor defense.
4.2.2 Proxy datasets $\mathcal{D}_{poison}$ and $\mathcal{D}_{clean}$
In this section, we introduce the process of extracting two proxy datasets from the suspected dataset $\mathcal{D}$: a proxy clean set $\mathcal{D}_{clean}$, which is identified with high probability as clean (i.e., they are likely to be clean though not guaranteed to be free from poisoning), and a proxy poison set $\mathcal{D}_{poison}$, which is identified with high confidence as being poisoned. Both proxy datasets are not required to be purely clean or poisoned. When the defender has no prior knowledge of the poisoned samples, various methods can be used to construct proxy datasets, such as sample diagnosis Gao et al. (2019); Huang et al. (2022); Guo et al. (2023) and trigger inversion Wang et al. (2019, 2022, 2023); Xu et al. (2024). If additional prior knowledge, such as a user-reported clean dataset, is available (as tested in Table 6), the performance of our method is expected to be improved. For simplicity, in this paper, we use random sampling to construct the proxy clean set $\mathcal{D}_{clean}$, based on the practical assumption that the poison samples will not dominate a training set. To construct the proxy poison set $\mathcal{D}_{poison}$, we adopt an existing backdoor sample detection method from SEEP He et al. (2024), which leverages an observation that backdoored models tend to exhibit over-confidence on poisoned samples Li et al. (2021b). SEEP utilizes training dynamics to identify anomalous samples in the training corpus, which are proven to have a high probability of being backdoor samples. In our example, we only focus on the most confident samples extracted (i.e., seed set). In Section 5.2, we show that extracting only 1% of the original poisoned data in $\mathcal{D}_{poison}$ and allowing even 20% poisoned data in $\mathcal{D}_{clean}$ are sufficient to launch an effective defense using our Greedy-Module-Substitution (GMS).
4.2.3 Greedy search for purified model $M_{pure}$
Using proxy datasets $\mathcal{D}_{poison}$ and $\mathcal{D}_{clean}$, we can estimate ASR and ACC to quantify the level of backdoor defense and task utility for a model $M$, we define the following two metrics:
$\delta_{\text{asr}}(M)$ $=s_{asr}(M_{v})-s_{asr}(M),$ (2)
$\delta_{\text{acc}}(M)$ $=s_{acc}(M_{v})-s_{acc}(M).$ (3)
Here, $s_{acc}(M)$ represents the CACC of $M$ on $\mathcal{D}_{clean}$, while $s_{asr}(M)$ denotes its ASR on $\mathcal{D}_{poison}$. Similarly, $s_{acc}(M_{v})$ and $s_{asr}(M_{v})$ are the ASR and CACC for the original victim model $M_{v}$. In this context, $\delta_{\text{asr}}(M)$ quantifies the reduction in ASR achieved by a purified model $M$, while $\delta_{\text{acc}}(M)$ reflects the drop in task utility. Plugging them back into the objective (Equation 1) gives us an score. [^2] An inaccurate $\mathcal{D}_{poison}$ may lead to a smaller $\delta_{\text{asr}}$, while an impure $\mathcal{D}_{clean}$ might result in a larger $\delta_{\text{acc}}(M)$. However, since the metrics are always calculated on fixed proxy datasets, these effects can be controlled by adjusting $\alpha$.
While our goal is to purify $M_{v}$ by replacing its modules across $L$ layers, an exhaustive search is computationally impractical due to its exponential complexity (e.g., $2^{6\cdot 12}$ for a BERT-base (Devlin et al., 2019)). Thus, we propose a greedy algorithm that transforms the problem into a feasible search by iteratively identifying and replacing the most critical module, maximizing the objective score to derive the optimal purified model $M_{pure}$.
Algorithm 1 Greedy-Module-Substitution (GMS)
1:  Input: Victim model $M_{v}$, proxy model $M_{p}$, module set $\mathcal{M}$, layer set $\mathcal{L}$, proxy datasets $\mathcal{D}_{clean}$ and $\mathcal{D}_{poison}$
2:  Output: Purified model $M_{pure}$
3:  Initialize $S\leftarrow\mathbf{1}_{|\mathcal{M}|\times|\mathcal{L}|}$, $c_{best}\leftarrow-\infty$, $S_{best}\leftarrow\varnothing$
4:  $(s_{acc},s_{asr})\leftarrow$ Evaluate($M_{v}$)
5:  # Iteratively Update Selected Modules and Layers:
6:  while $|\mathcal{M}|>1$ or $|\mathcal{L}|>1$ do
7:     $c^{*}\leftarrow-\infty$, $S^{*}\leftarrow\varnothing$
8:     for each $m\in\mathcal{M}$ or $l\in\mathcal{L}$ do
9:        $S^{\prime}\leftarrow S$ with $S^{\prime}_{:,m}=\mathbf{0}_{L\times 1}$ or $S^{\prime}_{l,:}=\mathbf{0}_{1\times|\mathcal{M}|}$
10:        $M_{pure}^{\prime}\leftarrow$ Replace($M_{v}$, $M_{p}$, $S^{\prime}$)
11:        $c\leftarrow$ ComputeScore($M_{pure},s_{acc},s_{asr}$)
12:        if $c>c^{*}$  then
13:           Update $c^{*}\leftarrow c,S^{*}\leftarrow S^{\prime}$
14:        end if
15:     end for
16:     $S\leftarrow S^{*}$ and update $\mathcal{M},\mathcal{L}$ accordingly
17:     if $c^{*}>c_{best}$  then
18:        Update $c_{best}\leftarrow c^{*}$, $S_{best}\leftarrow S^{*}$
19:     end if
20:     if $c_{best}$ not been updated for $T$ iterations then
21:        break
22:     end if
23:  end while
24:  return  $M_{pure}=\text{Replace}(M_{v},M_{p},S_{best})$
Purifying $M_{v}$ by module substitution
Given a backdoored model $M_{v}$ (characterized by both high CACC on a clean dataset and high ASR on a poisoned dataset), we aim to identify the specific modules and layers that, when substituted, allow the model to maintain strong performance on the clean task (i.e., minimizing $\delta_{\text{acc}}$) while eliminating the backdoor features (i.e., maximizing $\delta_{\text{asr}}$).
Our algorithm, as demonstrated in Figure 2 and Algorithm 1, tracks the parameters selected for substitution using a module set ($\mathcal{M}$) and a layer set ($\mathcal{L}$), e.g., if =$\mathcal{M}=\{O,F,P\}$ and =$\mathcal{L}=\{7,8,9\}$, it indicates that the $O$, $F$, and $P$ module parameters in the 7th, 8th, and 9th layers of the victim model $M_{v}$ should be replaced with the corresponding parameters from proxy model $M_{p}$. For simplicity, we introduce a substitution matrix $S\in\{0,1\}^{|\mathcal{M}|\times|\mathcal{L}|}$ to specify which modules in which layer in the victim model $M_{v}$ are selected for substitution. Note that while the shape of $S$ is determined by the initial sizes of sets $\mathcal{M}$ and $\mathcal{L}$, these two sets are iteratively updated throughout the algorithm. Each row of $S$ corresponds to a layer in the layer set $\mathcal{L}$, and each column corresponds to a module in the module set $\mathcal{M}$. The value of $S[l,m]=1$ indicates that the module $m\in\mathcal{M}$ in layer $l\in\mathcal{L}$ in the victim model $M_{v}$ is selected for substitution, and $S[l,m]=0$ otherwise. Given $S$, we can succinctly represent the parameter substitution as $S\odot M_{p}+(1-S)\odot M_{v}$, where $\odot$ denotes the element-wise product.
We start by selecting all modules in $M_{v}$ to be replaced with those of $M_{p}$, i.e., =$\mathcal{M}=\{K,Q,V,O,F,P\}$, =$\mathcal{L}=\{1,\ldots,L\}$, $S=\mathbf{1}_{|\mathcal{M}|\times|\mathcal{L}|}$. At $t$-th iteration, the algorithm considers two types of updates to $S^{t}$: deselecting a module from the current module set $\mathcal{M}$ (affecting all selected layers) or deselecting a layer from the layer set $\mathcal{L}$ (affecting all selected modules). The update that maximizes the score in Equation 1 is applied to $S^{t+1}$. The process continues until one of two conditions is met: (1) further iterations will not yield a better strategy (controlled by stopping patience $T$); or (2) no additional layers or modules can be removed, i.e., $|\mathcal{M}|=|\mathcal{L}|=1$. Finally,the strategy $S$ that yields a purified model $M_{pure}$ with the highest score is returned. The complexity of our algorithm is a practical quadratic $O(|\mathcal{M}|^{2}+|\mathcal{L}|^{2})$.
5 Experiments
5.1 Experimental settings
Datasets.
We evaluate our method on four datasets: SST-2 (Socher et al., 2013), OLID (Zampieri et al., 2019), MNLI (Williams et al., 2018), and AGNEWS (Zhang et al., 2015). These datasets cover text classification and natural language inference (NLI) tasks, including binary and multi-class classification scenarios, and are widely used for evaluating text backdoor attacks and defenses (Qi et al., 2021a; He et al., 2023; Gupta and Krishna, 2023; Arora et al., 2024).
We download the open-source versions of these datasets provided by HuggingFace (Lhoest et al., 2021). Table 1 presents the statistics of the evaluated datasets. To train backdoored models, we construct poisoned datasets by injecting triggers into the training split, while the validation split is used to construct a clean and a poison test set respectively. For MNLI, we use a random subset of 100,000 samples to reduce overhead.
Backdoor Attacks.
We study defenses against four prominent types of data-poisoning backdoor attacks: (1) BadNets (Kurita et al., 2020), (2) InsertSent (Dai et al., 2019), (3) Learnable Word Substitution (LWS(Qi et al., 2021d), and (4) HiddenKiller (Qi et al., 2021c). The first two correspond to insertion-based methods, while the latter two involve synonym substitution and syntactic paraphrasing approaches, respectively.
Follow Arora et al. (2024), we use rare words {"cf", "mn", "bb", "tq", "mb" } as triggers for BadNets and phrases {"I watched this movie", "no cross, no crown" } for InsertSent. The poison target labels for each dataset are listed in Table 1. While our main experiments focus on 20% poison rate (in line with the literature (Dai et al., 2019; Qi et al., 2021c; Arora et al., 2024)), i.e., 20% of training samples are poisoned, we also evaluate lower poison rates of 10%, 5% in Section B.2.
Dataset Classes Train Test Target Class
Clean Poison
SST-2 2 67,349 872 444 Negative (0)
OLID 2 13,240 860 240 Not offensive (1)
MNLI 3 100,000 400 285 Neutral (1)
AGNews 4 120,000 7,600 5,700 Sports (1)
Table 1: The statistics of our evaluated datasets.
Defense Baselines.
We choose six renowned defenses as our baselines, including two data-wise detection methods: (1) ONION (Qi et al., 2021a) and (2) Z-Def. (He et al., 2023), and four model-wise purification methods: (3) PURE (Zhao et al., 2024), (4) TIES (Yadav et al., 2023), (5) DARE (Yu et al., 2024), and (6) WAG (Arora et al., 2024).
Among these, ONION and Z-Def. detect and remove triggers from datasets. ONION identifies outlier words (potential triggers) using language models, e.g., GPT-2 (Radford et al., 2019), while Z-Def. detects spurious correlations between tokens and labels. In contrast, (2)-(6) are model purification methods. PURE purifies the victim model using attention head pruning and normalization techniques. TIES, DARE and WAG merge the victim model with other models to sanitize it.
Evaluation Metrics.
As introduced in  Section 3, we adopt Clean Accuracy (CACC) and Attack Success Rate (ASR) to measure the model performance. CACC is evaluated on a clean test set, where both the samples and labels are ground truth. In contrast, ASR is evaluated on a poisoned test set, where triggers are implanted into each sample, and the labels are flipped to the target class.
BadNet InsertSent LWS HiddenKiller
[…]
Benign 4.1 95.9 2.2 95.9 12.8 95.9 16.5 95.9
Victim 100.0 96.0 100.0 96.3 98.0 95.4 96.5 95.7
Proxy Model (IMDB) 7.4 89.1 4.3 89.1 10.8 89.1 13.7 89.1
ONION 56.8 92.9 99.9 93.3 85.7 91.9 92.9 92.9
Z-Def. 4.6 96.1 1.8 95.6 97.3 95.3 35.7 95.4
PURE 0 50.9 0 50.9 0 50.9 0 50.9
TIES 99.9 95.7 100.0 95.8 93.5 95.2 88.8 95.3
DARE w/ TIES 99.3 96.0 100.0 96.2 96.4 95.7 92.7 95.8
WAG 84.4 94.8 60.1 95.2 58.8 94.8 56.2 92.5
SST-2 Ours (GMS) 4.5 91.6 1.9 92.5 9.7 91.7 10.4 91.2
Benign 1.9 95.4 0.5 95.4 0.5 95.4 1.1 95.4
Victim 99.9 95.1 99.6 95.3 99.6 94.5 100.0 95.1
Proxy Model (BBCNews) 1.5 70.2 1.7 70.2 1.8 70.2 3.4 70.2
ONION 59.4 94.8 97.8 95.1 84.8 94.5 99.6 94.7
Z-Def. 1.6 95.3 0.4 95.3 97.9 96.1 100.0 95.0
PURE 2.8 86.3 3.0 85.4 6.5 85.0 9.4 84.6
TIES 99.9 94.6 99.6 94.4 97.7 95.8 100.0 94.4
DARE w/ TIES 99.9 95.2 99.6 95.4 97.8 96.5 100.0 95.2
WAG 92.7 94.1 97.8 94.4 78.0 93.9 90.9 94.3
AGNews Ours (GMS) 2.5 91.0 2.4 92.6 3.2 91.7 6.5 90.4
Table 2: (Partial) Performance of our method on two datasets compared to baselines under various backdoor attacks on the RoBERTa-large model, with each value averaged over three seeds. We highlight the top-2 lowest ASR results in blue cells, and the highest CACC results in bold. Results for other datasets are in Table 8.
Implementation details
Following Arora et al. (2024), we compare all methods on RoBERTa-large (Liu et al., 2019) and BERT-base-uncased (Devlin et al., 2019). All victim models are fine-tuned on poisoned datasets for three epochs using the Adam optimizer with no weight decay (Kingma and Ba, 2015) and a learning rate of $2\times 10^{-5}$. All defense baselines are implemented based on their respective open-source repositories (see Section A.3). The high-confidence proxy datasets ($\mathcal{D}_{clean}$ and $\mathcal{D}_{poison}$) are extracted following Section 4.2.2. To train proxy models for selection, we use IMDB (Zhang et al., 2015) for the SST-2 victim model, Twitter Abusive (Founta et al., 2018) for the OLID, SNLI (Young et al., 2014) for MNLI, and BBCNews (Greene and Cunningham, 2006) for AGNews.
Our method uses hyperparameter $\alpha$ in Equation 1 to control the trade-off between backdoor defense and task utility. By default, $\alpha$ is set to 0.4 but can be adjusted based on priorities: smaller values (e.g., 0.1) for lower ASR and larger values (e.g., 1.0) for higher utility. The stopping patience $T=5$ is default in Algorithm 1. In Section B.3, we justify these choices and show that our method is NOT particularly sensitive to hyperparameters.
5.2 Main results
Table 2 presents the performance of our method compared to all baselines under various backdoor attacks on the RoBERTa-large model, with scores averaged over three runs using different seeds. Results on other datasets and architectures are provided in Section B.1. Across all datasets and backdoor attacks, our method consistently ranks among the top 2, regarding backdoor removal performance, with minor harm to clean accuracy. In particular, under the two particularly challenging backdoor tasks, LWS and HiddenKiller, GMS demonstrates significant improvements (i.e., at least 25%) over all baselines—for example, for SST-2, our method reduces the ASR on LWS to 9.7%, compared to 58.8% for the following baseline Z-Def. While Z-Def achieves competitive performance with GMS in defending against BadNet and InsertSent, Z-Def is much less effective against LWS on both datasets. This is because Z-Def relies on lexical and syntactic features to detect outliers in the poisoned dataset, whereas LWS attacks subtly replace words with synonyms, effectively bypassing outlier detection. As for PURE, another competitive recent approach, we found its head-pruning step is highly unstable across both tasks and architectures. For example, it performs well on the BERT-base model in Table 9 (following the settings reported in PURE (Zhao et al., 2024)), but performs poorly on most tasks with RoBERTa-large. We discuss the potential reasons for this instability in Section B.1.
Dataset Model CACC $\uparrow$
SST-2 Benign Victim Model 95.9
Proxy Model (IMDB) 88.7
Purified Model (GMS) 95.8
AGNews Benign Victim Model 95.2
Proxy Model (BBCNews) 78.1
Purified Model (GMS) 95.3
Table 3: CACC of benign victim models preserved.
BadNet InsertSent LWS HiddenKiller
Dataset Proxy Model ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$
IMDB 5.4 92.8 1.6 94.0 12.6 93.8 11.0 93.0
YELP 4.3 96.8 8.1 94.8 23.2 92.3 19.1 93.8
SST-2 AMAZON 3.8 92.9 1.1 95.8 20.1 92.6 13.5 93.8
Table 4: Proxy models trained on different datasets can all effectively purify victim models (trained on the SST-2).
Model Dataset Backdoored Method Before Substitution After Substitution
[…]
Victim SST-2 BadNet 95.6 100.0 95.3 4.5
Proxy SST-2 HiddenKiller 95.4 7.9
Victim SST-2 HiddenKiller 95.4 96.7 95.9 14.0
Proxy SST-2 BadNet 95.6 17.1
Victim SST-2 BadNet (Different triggers) 95.6 100.0 87.8 6.8
Proxy IMDB 87.7 8.6
Victim SST-2 BadNet (Same triggers) 95.6 100.0 95.8 100.0
Proxy IMDB 90.1 100.0
Table 5: Proxy models with implanted backdoors can still effectively purify victim models if the backdoor does not exactly match in both attack strategy and trigger.
Additionally, although merging multiple models has been shown to be an effective backdoor defense, we observed that applying all merging baselines using a single proxy model is sub-optimal, aligning with the results in Arora et al. (2024). We found these methods prioritize preserving accuracy over removing backdoors. This behavior likely arises from the design of merging mechanisms, which are intended to preserve the performance of each merged model on downstream tasks. Consequently, backdoor tasks are treated equivalently to downstream tasks, exposing a persistent vulnerability.
5.3 Ablation studies
Utility Preservation for Benign Victim Models
While all our previous experiments focus on malicious victim models, it is also important to consider scenarios where the victim model is benign. Does our method compromise the utility of a clean victim model? As shown in Table 3, our results effectively dispel this concern. When applying GMS to a benign model, the CACC remains unaffected.
The effect of different proxy models
Our method relies on a proxy model for substitution, raising the natural question of how sensitive it is to the selection of proxy models. We tested two scenarios: Scenario 1: We evaluated the performance of using different proxy models trained on homologous datasets, such as IMDB, Yelp, and Amazon (Zhang et al., 2015), to purify a victim model trained on SST-2. As shown in Table 4, our method is largely insensitive to the choice of proxy model across all backdoor attacks. Using any proxy model, GMS effectively defends against all tested attacks. Scenario 2: What if the proxy model itself is backdoored? We tested cases where the victim model and the proxy model were trained on the same dataset with different backdoor attacks or on different datasets with the same attacks. From Table 5, we observed that as long as the backdoor in the proxy model is not identical to that in the victim model (i.e., the same backdoor strategy with the same backdoor trigger), our GMS method consistently removes the backdoor.
Performance under different knowledge priors
While our method uses random sampling and outlier detection to extract the proxy datasets $\mathcal{D}_{clean}$ and $\mathcal{D}_{poison}$, we also evaluate the potential performance gain if more accurate prior knowledge were available. For example, we may receive an accurate clean dataset or an accurate label-flipped poison dataset (i.e., dirty label) from reliable users or annotators. As shown in Table 6, with or without prior knowledge of accurate clean or poisoned data, our method consistently purifies the backdoor, achieving an ASR same as that of a benign model (4.1%). With optimal priors, GMS maintains the CACC of the benign model (95.6%), compared to 93.6% in the worst case without such prior knowledge.
Knowledge Prior Merged SST-2 (BadNet) & IMDB
[…]
Random Sampling Pure Clean Outlier Detection Dirty Label
3.2 93.6
4.1 94.9
4.3 92.7
4.3 95.6
Benign SST-2 4.1 95.6
Table 6: Performance of purifying a victim model trained on SST-2 with a proxy model trained on IMDB, using proxy datasets with varying knowledge priors. The backdoor attack strategy is BadNet.
BadNet InsertSent LWS HiddenKiller
[…]
Benign 4.1 95.9 2.2 95.9 12.8 95.9 16.5 95.9
Victim 100.0 96.0 100.0 96.3 98.0 95.4 96.5 95.7
Proxy Model (IMDB) 7.4 89.1 4.3 89.1 10.8 89.1 13.7 89.1
GMS 4.5 91.6 1.9 92.5 9.7 91.7 10.4 91.2
SST-2 SST-2 BadNet Strategy 4.5 91.6 9.8 93.0 22.3 92.8 13.2 90.3
Benign 1.9 95.4 0.5 95.4 0.5 95.4 1.1 95.4
Victim 99.9 95.1 99.6 95.3 99.6 94.5 100.0 95.1
Proxy Model (BBCNews) 1.5 70.2 1.7 70.2 1.8 70.2 3.4 70.2
GMS 2.5 91.0 2.4 92.6 3.2 91.7 6.5 90.4
AGNews AGNews BadNet Strategy 1.4 90.7 1.2 89.8 2.6 90.1 12.0 90.2
Table 7: (Partial) The performance of transferring strategy searched based on BadNet to other attacks. Results for other datasets are in Table 12.
Transferability of substitution strategies
Interestingly, upon examining the module substitution strategies ($S_{best}$) searched by GMS for addressing various backdoor attacks (illustrated with examples in Figure 6 in Section B.5), we observe that, despite differences among the strategies, they exhibit transferability across different backdoor attacks.
In Table 7, for each task dataset, we apply the substitution strategy found on BadNet (replacing the fewest modules and layers as shown in Figure 6) to to determine which modules to substitute in the three other victim models under different attacks, replacing them with modules from their corresponding proxy models, i.e., directly executing Line 24 in Algorithm 1 using $S_{badnet}$. Compared with the complete GMS (i.e., searching for the optimal strategies $S_{best}$ for each attack), the transferred strategy consistently performs comparably. This demonstrates that for defenders who only have access to the victim model and are unaware of the victim dataset, a universal GMS defense strategy exists for each task that can effectively defend against multiple attacks.
6 Conclusion
We propose Greedy-Module-Substitution (GMS), an efficient and effective model purification defense against data-poisoning attacks in NLP, eliminating the need for strictly clean datasets or clean auxiliary models. GMS recursively identifies modules critical to backdoor pathways and substitutes them for those from a proxy model, which is not necessarily to be clean. Extensive experiments demonstrate the effectiveness of GMS and, more excitingly, the transferability of the optimized substitution strategies.
Limitations
Our method goes one step beyond previous auxiliary-model-based defenses by allowing the proxy model to be malicious, as long as its backdoor is not identical to the victim’s backdoor (i.e., the same attack method with exactly the same trigger, which is rare in practice) as shown in Table 5. However, further relaxation of the requirements to handle cases where the proxy model is backdoored in the same way could enable even more robust and powerful defense strategies. Exploring this direction presents an important and meaningful avenue for future research.
Appendix A Experiment details
A.1 Experiment setup
We adopt the same training configuration as WAG (Arora et al., 2024) to ensure a fair comparison. This includes fine-tuning roberta-large and bert-base-uncased on poisoned datasets for three epochs, using the Adam optimizer with no weight decay (Kingma and Ba, 2015) and a learning rate of $2\times 10^{-5}$. The batch size and maximum sequence length are set to 32 and 128, respectively.
To validate our defense, we first adapt the aforementioned backdoor methods and datasets to obtain victim models. Next, we train proxy models on homologous datasets to prepare for future merging. We then construct high-confidence proxy-clean ($\mathcal{D}_{clean}$) and proxy-poison ($\mathcal{D}_{poison}$) sets from the training data (using the methods illustrated in Section 4.2.2) to guide our parameter substitution strategy search.
For a model trained on unverified data, access to fully clean data is often not guaranteed. However, the corresponding task is typically known to the defender, providing an opportunity to identify homologous models trained on overlapping tasks that can be used for merging. For instance, previous works, WAG and PURE, utilized the IMDB (Maas et al., 2011) dataset to purify models trained on poisoned SST-2 datasets, as both share the sentiment classification task domain.
In this work, we select proxy models that share a similar domain with the poisoned tasks. Specifically, we use Twitter Abusive (Founta et al., 2018) for OLID, SNLI [^3] Creative Commons Attribution-ShareAlike 4.0 International License (Young et al., 2014) for MNLI, and BBCNews (Greene and Cunningham, 2006) for AGNews. To evaluate generalizability, we use three sentiment classification datasets—IMDB, YELP, and Amazon Reviews (Zhang et al., 2015)—to purify victim SST-2 models through model merging. All datasets are downloaded from Hugging Face, which adheres to the Apache License 2.0.
Our method involves only one hyperparameter, $\alpha$, as described in  Equation 1, which controls the preference between model utility and attack resistance during the model merging strategy search. We set $\alpha$ to 0.4 by default, which slightly favors seeking a more attack-resistant model. It can be adjusted to smaller values (e.g., 0.1) when a lower ASR is the primary target, and to larger values (e.g., 1.0) when high utility is demanded.
A.2 Computational Resources
We conduct experiments using three seeds on Nvidia V100 and A1000 GPUs, and report the average scores. Running our method takes only 4 minutes on a single GPU.
BadNet InsertSent LWS HiddenKiller
[…]
Benign 4.1 95.9 2.2 95.9 12.8 95.9 16.5 95.9
Victim 100.0 96.0 100.0 96.3 98.0 95.4 96.5 95.7
Proxy Model (IMDB) 7.4 89.1 4.3 89.1 10.8 89.1 13.7 89.1
ONION 56.8 92.9 99.9 93.3 85.7 91.9 92.9 92.9
Z-Def. 4.6 96.1 1.8 95.6 97.3 95.3 35.7 95.4
PURE 0 50.9 0 50.9 0 50.9 0 50.9
TIES 99.9 95.7 100.0 95.8 93.5 95.2 88.8 95.3
DARE w/ TIES 99.3 96.0 100.0 96.2 96.4 95.7 92.7 95.8
WAG 84.4 94.8 60.1 95.2 58.8 94.8 56.2 92.5
SST-2 Ours (GMS) 4.5 91.6 1.9 92.5 9.7 91.7 10.4 91.2
Benign 29.0 84.9 31.0 84.9 45.6 84.9 53.6 84.9
Victim 99.9 85.2 100.0 85.0 94.5 85.1 100.0 85.0
Proxy Model (Twitter) 37.8 84.4 40.0 84.4 55.4 84.4 67.1 84.4
ONION 75.0 84.8 99.4 84.8 86.1 84.1 99.6 84.7
Z-Def. 29.7 85.6 30.3 85.3 93.5 85.3 53.9 85.7
PURE 82.2 75.2 87.5 75.2 76.1 79.1 100.0 72.1
TIES 36.2 84.8 38.2 84.8 57.5 84.5 65.3 84.5
DARE w/ TIES 95.6 86.3 77.8 86.0 90.5 85.9 86.8 86.3
WAG 52.0 85.0 48.6 84.6 63.1 84.5 68.2 85.0
OLID Ours (GMS) 32.1 85.0 28.0 84.2 52.2 84.4 64.2 84.9
Benign 12.3 87.6 12.6 87.6 26.4 87.6 36.9 87.6
Victim 100.0 89.4 100.0 90.3 96.0 89.0 99.9 89.4
Proxy Model (SNLI) 12.2 84.1 9.2 84.1 25.3 84.1 31.7 84.1
ONION 64.3 86.1 98.6 86.9 89.0 85.5 98.8 86.6
Z-Def. 11.1 88.3 11.6 89.7 92.2 89.1 50.6 89.7
PURE 33.3 33.8 33.3 33.8 33.3 33.8 33.3 33.8
TIES 91.6 89.2 94.5 89.5 80.7 89.2 88.6 90.0
DARE w/ TIES 93.5 90.5 100.0 91.4 93.9 89.9 99.4 90.5
WAG 71.0 88.7 60.3 88.5 77.9 88.0 80.5 88.8
MNLI Ours (GMS) 10.8 86.5 10.7 86.3 14.0 86.5 31.7 86.3
Benign 1.9 95.4 0.5 95.4 0.5 95.4 1.1 95.4
Victim 99.9 95.1 99.6 95.3 99.6 94.5 100.0 95.1
Proxy Model (BBCNews) 1.5 70.2 1.7 70.2 1.8 70.2 3.4 70.2
ONION 59.4 94.8 97.8 95.1 84.8 94.5 99.6 94.7
Z-Def. 1.6 95.3 0.4 95.3 97.9 96.1 100.0 95.0
PURE 2.8 86.3 3.0 85.4 6.5 85.0 9.4 84.6
TIES 99.9 94.6 99.6 94.4 97.7 95.8 100.0 94.4
DARE w/ TIES 99.9 95.2 99.6 95.4 97.8 96.5 100.0 95.2
WAG 92.7 94.1 97.8 94.4 78.0 93.9 90.9 94.3
AGNews Ours (GMS) 2.5 91.0 2.4 92.6 3.2 91.7 6.5 90.4
Table 8: Performance of our method compared to baselines under various backdoor attacks on the RoBERTa-large model, with each value averaged over three seeds.
A.3 Implementation details for baselines
We follow the open-source implementations for each baseline, and basically using their default hyper-parameters, while maintaining using the identical datasets, backdoor settings and the trained models for a fair comparison.
For the baseline DARE Yu et al. (2024), it first applies random parameter dropping and rescaling to the involved models with a specified drop rate, and then incorporates model merging techniques to combine the processed models. Various model merging methods can be integrated with DARE, and we choose TIES merging as a representative, as it demonstrates decent performance for encoder-based models (e.g., bert-base, roberta-base). Their method involves three tunable hyperparameters for encoder-based LMs, as outlined in Table 5 of their paper: drop rate, scaling term, and ratio to retain. We retained the original search space for drop rate and scaling term but expanded the ratio to retain from [0.1, 0.2, 0.3] to [0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8] for the roberta-large model to ensure more complete search.
For the baseline PURE (Zhao et al., 2024), it first trains the victim model on a proxy clean dataset (from the same domain or a similar task, e.g., poisoned on IMDB and fine-tuned on SST-2) and prunes the attention heads that cause attention drifting between poisoned and clean texts. The pruned model is then fine-tuned further on the proxy clean dataset to normalize the remaining attention heads for purification. To ensure a fair comparison in the main experiments with the roberta-large model, we use the same poisoned and proxy datasets, e.g., poisoned on the SST-2 dataset with IMDB as the proxy clean dataset. We set the hyperparameter accuracy threshold (used to stop head pruning) to 0.90 for SST-2, MNLI, and AGNews, and 0.85 for OLID to prevent overly aggressive pruning. For the bert-base model, we follow the original implementation, including the use of the SST-2 dataset for fine-tuning SST-2 victim models and maintaining the default accuracy threshold of 0.85. While PURE uses the label flip rate (LFR) as its evaluation metric for backdoor defense (implanting triggers into test data while keeping the labels unchanged), we adopt the attack success rate (ASR) on label-flipped test data as our evaluation metric for a fair comparison with our method.
We adhere to the default implementations and hyperparameter settings for all other baseline methods.
Appendix B Additional results
BadNet InsertSent LWS HiddenKiller
[…]
Benign 8.5 92.9 3.5 92.9 22.2 92.9 17.3 92.9
Victim 100.0 92.9 100.0 92.2 98.1 91.5 95.9 91.8
Proxy Model (IMDB) 10.4 85.4 6.4 85.4 17.9 85.4 14.7 85.4
ONION 58.0 89.9 99.7 89.9 85.4 88.6 94.4 89.2
Z-Def. 8.3 92.5 1.8 91.9 97.4 90.9 38.7 91.5
PURE 30.1 92.0 10.7 91.5 66.4 91.1 28.7 91.3
TIES 99.2 92.7 98.2 92.4 91.2 92.3 86.8 92.6
DARE w/ TIES 100.0 93.0 90.8 88.9 92.6 92.7 94.1 90.6
WAG 74.3 91.9 70.7 91.7 73.5 91.7 60.3 91.6
SST-2 Ours (GMS) 10.8 86.8 9.8 88.8 27.4 88.0 19.5 86.8
Table 9: Performance of our method compared to baselines under various backdoor attacks on the BERT-base model, with each value averaged over three seeds. The victim model is trained on SST-2 dataset.
B.1 Performance across different model architectures and different dataset
Performance across different dataset. Due to space constraints, we present the complete performance results of our method compared to all baselines under various backdoor attacks on the RoBERTa-large model across all four datasets in Table 2. The settings are consistent with those in Section 5.2, with each score averaged over three runs using different seeds. For the OLID dataset, we specifically set $\alpha=0.1$ to enable more aggressive defense, as shown in Figure 4. The top two ASR performances are highlighted. Across all datasets and backdoor attacks, our method consistently ranks among the top two in backdoor removal performance, with minimal harm to clean accuracy. Notably, under the two particularly challenging backdoor tasks, LWS and HiddenKiller, GMS achieves significant improvements (at least 25%) over all baselines. For instance, on SST-2, GMS reduces the ASR for LWS to 9.7%, compared to 58.8% for the next-best baseline, Z-Def.
While Z-Def achieves competitive performance with GMS in defending against BadNet and InsertSent, Z-Def is much less effective against LWS on both datasets. This is because Z-Def relies on lexical and syntactic features to detect outliers in the poisoned dataset, whereas LWS attacks subtly replace words with synonyms, effectively bypassing outlier detection. As for PURE, another competitive recent approach, we found its head-pruning step is highly unstable across both tasks and architectures. For example, it performs well on the BERT-base model in Table 9 (following the settings reported in PURE (Zhao et al., 2024)), but performs poorly on most tasks with RoBERTa-large. This instability seems to arise from PURE’s reliance on accuracy from the clean proxy dataset as the pruning stopping criterion. For well-trained models with high accuracy, a substantial number of heads are pruned, leading to a broken purified model (e.g., 0% ASR but random-guess-level CACC).
Performance across different model architectures. We also evaluated our method on different architectures. Table 9 shows the results on the BERT-base model. The poison rate remains at 20%, and each model is trained for three epochs. Scores are reported under $\alpha=0.4$ and averaged over three seeds: 1000, 2000, and 3000. Across all attacks, our method performs consistently well, particularly excelling against more challenging attack strategies. Notably, it achieves over a 39% advantage in defending against LWS and a 9% improvement over the second-best baseline for HiddenKiller.
B.2 Performance under different poison rate: 20%, 10% and 5%
While the experiments in the main paper are conducted with a 20% poison rate to ensure fair comparison with the baselines, following the settings in (Dai et al., 2019; Qi et al., 2021c; Arora et al., 2024), i.e., 20% of the training samples are poisoned, we also evaluate lower poison rates of 10% and 5%, as shown in Table 10. Our method is effective in model purification across all poison rates.
Victim Proxy Poison Rate Metric Result
[…]
Table 10: Results for SST-2 with varying poison rates.
Figure 3: Results of using different weights (alpha) on OLID dataset.
Figure 4: Results of using different weights (alpha) on AGNEWS dataset.
B.3 How sensitive is our method to the selection of hyperparameters?
Our method involves two hyperparameters: $\alpha$ in Equation 1 and stopping patience $T$, with default values of $\alpha=0.4$ and $T=5$ used in main results. Below, we justify these choices. As shown in Figure 4 and Figure 4, for most dataset and backdoor attack combinations, the default value $\alpha=0.4$ strikes a balanced trade-off between utility and backdoor defense strength. An exception is the HiddenKiller attack on the OLID dataset, where a more aggressive $\alpha=0.1$ is recommended. This is because OLID is a dataset collected from tweets, while HiddenKiller paraphrases the data using formal syntactic templates, drastically altering OLID’s language style. This change significantly impacts CACC, necessitating a higher weight for backdoor removal to ensure effective purification. For stopping patience, as shown in Figure 5, the score computed in Equation 1 generally follows a (weakly) monotonically decreasing trend. While a larger $T$ improves performance, we choose $T=5$ as a reasonable balance between performance and efficiency.
Figure 5: The search iteration history for four kinds of backdoors on SST-2 dataset.
Methods
Metrics Stage WAG PURE Clean Ours
Before 94.72 93.12 95.76 91.86
[…]
Before 84.40   6.31   4.05   3.83
[…]
Table 11: Evaluating the robustness of different parameter purification methods towards the Retuning Attack (RA) using 0.1% of the original poisoned training set (which contains 20% poisoned data) for fine-tuning. We report the performance of each methods before and after the RA.
Figure 6: The optimal substitution strategy for defending against each backdoor attack for the roberta-large model trained on the SST-2 dataset. The green squares indicate the substituted modules.
B.4 Robustness to Retuning Attacks
Recent works Min et al. (2024); Qi et al. (2024) have highlighted a critical pitfall in backdoor defenses: purified models with low ASR often fail to completely eliminate inserted backdoor features. A straightforward Retuning Attack (RA)—which involves fine-tuning a purified model on a small number of backdoored samples (e.g., 0.2% of the original poisoned data) mixed with benign samples (to maintain clean Accuracy on the main task)—can easily recover the backdoor in the purified model, as demonstrated in Table 11.
Several related studies Qi et al. (2024); Chen et al. (2024b); Tarun et al. (2024) have sought to uncover the underlying reasons for this failure. A key observation is that the purified model’s parameters do not sufficiently deviate from those of the backdoored model. Specifically, the weight differences between a purified model and a (retrained) benign model are significantly larger than those between the purified and backdoored models. This suggests that purified and backdoored models are connected via a backdoor-related path in the loss landscape Min et al. (2024). Such vulnerabilities pose severe risks, as purified models are often deployed in various downstream applications, and even with backdoor defenses applied, attackers can easily re-trigger the backdoor in downstream tasks Min et al. (2024).
This challenge motivates us to propose a purification method aimed at breaking this backdoor-connected path. Intuitively, achieving this requires completely replacing "suspected" parameters inherited from the backdoored model, rather than pruning (Zhao et al., 2024) or merging them with other parameters Arora et al. (2024); Yadav et al. (2023). Moreover, because parameters often interact within functional modules, modifying at the module level rather than the individual parameter level tends to remove backdoor features more effectively. As shown in Table 11, when using 0.1% of the original poisoned training set to launch the Retuning Attack (RA) Min et al. (2024), the parameter-merging baseline, WAG Arora et al. (2024), is easily re-triggered. In contrast, our module-substitution method demonstrates greater robustness while maintaining utility, performing comparably to the clean model baseline (i.e., finetuning a clean model on the retuning dataset). We provide our baseline PURE with additional advantages by granting access to the full clean dataset of the victim task for finetuning, as obtaining a usable purified model otherwise proves challenging. While the recent method PURE also demonstrates competitive performance, it is worth noting that significant effort was required to find a configuration where its pruning step did not break the purified model. Even with these adjustments, our method still substantially outperforms PURE, achieving an ASR of 11.93% compared to 19.37%—nearly twice the robustness.
B.5 Examples of Searched Substitution Strategy
We applied our GMS defense to purify victim models compromised by various backdoor attacks across four datasets. Examples of the searched strategies from the SST-2 dataset are shown in Figure 6. While the obtained strategies vary, we speculate that they are transferable, as adapting the minimum replacement strategy identified for BadNet to the other three attacks yields effective defense results, as discussed in Section 5.3 and  Section B.6.
BadNet InsertSent LWS HiddenKiller
[…]
Benign 4.1 95.9 2.2 95.9 12.8 95.9 16.5 95.9
Victim 100.0 96.0 100.0 96.3 98.0 95.4 96.5 95.7
Proxy Model (IMDB) 7.4 89.1 4.3 89.1 10.8 89.1 13.7 89.1
GMS 4.5 91.6 1.9 92.5 9.7 91.7 10.4 91.2
SST-2 SST-2 BadNet Strategy 4.5 91.6 9.8 93.0 22.3 92.8 13.2 90.3
Benign 29.0 84.9 31.0 84.9 45.6 84.9 53.6 84.9
Victim 99.9 85.2 100.0 85.0 94.5 85.1 100.0 85.0
Proxy Model (Twitter) 37.8 84.4 40.0 84.4 55.4 84.4 67.1 84.4
GMS 32.1 85.0 28.0 84.2 52.2 84.4 64.2 84.9
OLID OLID BadNet Strategy 33.6 85.0 37.4 84.6 56.5 84.3 65.0 84.7
Benign 12.3 87.6 12.6 87.6 26.4 87.6 36.9 87.6
Victim 100.0 89.4 100.0 90.3 96.0 89.0 99.9 89.4
Proxy Model (SNLI) 12.2 84.1 9.2 84.1 25.3 84.1 31.7 84.1
GMS 10.8 86.5 10.7 86.3 14.0 86.5 31.7 86.3
MNLI MNLI BadNet Strategy 12.2 86.3 10.8 86.3 27.0 85.2 35.2 86.4
Benign 1.9 95.4 0.5 95.4 0.5 95.4 1.1 95.4
Victim 99.9 95.1 99.6 95.3 99.6 94.5 100.0 95.1
Proxy Model (BBCNews) 1.5 70.2 1.7 70.2 1.8 70.2 3.4 70.2
GMS 2.5 91.0 2.4 92.6 3.2 91.7 6.5 90.4
AGNews AGNews BadNet Strategy 1.4 90.7 1.2 89.8 2.6 90.1 12.0 90.2
Table 12: The performance of transferring strategy searched based on BadNet to other attacks.
B.6 Transferability of substitution strategies
We present the results for SST-2 and AGNews in Table 7, with additional results for other datasets provided in Table 12. All experiments follow a consistent setup: the substitution strategy is optimized based on BadNet attacks and then applied to coordinate module substitutions for other attacks. For instance, in the MNLI task, the substitution strategy identified for defending against BadNet—using a proxy model trained on SNLI—is directly applied to replace modules in InsertSent, LWS, and HiddenKiller with those from the proxy model. Across all tasks, the defense performance is surprisingly strong, revealing an intriguing observation: a substantial overlap exists in the components impacted by different attacks.