### v2

# Cut the Deadwood Out: Backdoor Purification via Guided Module Substitution

Abstract
Model NLP models are commonly trained (or fine-tuned) on datasets from untrusted platforms like HuggingFace, posing significant risks of data poisoning attacks. A practical yet underexplored challenge arises when such backdoors are discovered after model deployment, making retraining-required defenses less desirable due to computational costs and data constraints. In this work, we propose Guided Module Substitution (GMS), an effective retraining-free method based on guided merging of the victim model with just a single proxy model. Unlike prior ad-hoc merging defenses, GMS uses a guided trade-off signal between utility and backdoor to selectively replaces modules in the victim model. GMS offers four desirable properties: (1) robustness to the choice and trustworthiness of the proxy model, (2) applicability under inaccurate data knowledge, (3) stability across hyperparameters, and (4) transferability across different attacks. Extensive experiments on encoder models and decoder LLMs demonstrate the strong effectiveness of GMS. GMS significantly outperforms even the strongest defense baseline, particularly against challenging attacks like LWS. The code is available at https://github.com/weijun-l/guided-module-substitution.
Cut the Deadwood Out: Backdoor Purification
via Guided Module Substitution
Yao Tong1thanks: Equal contribution., Weijun Li2∗, Xuanli He3, Haolan Zhan4, Qiongkai Xu2thanks: The corresponding author. 1 National University of Singapore, 2 Macquarie University 3 University College London, 4 Monash University tongyao@u.nus.edu, weijun.li1@hdr.mq.edu.au xuanli.he@ucl.ac.uk, haolan.zhan@monash.edu, qiongkai.xu@mq.edu.au
1 Introduction
Modern NLP models are frequently adapted to specific downstream tasks through fine-tuning on custom datasets Howard and Ruder (2018). In practice, these datasets are often collected from diverse sources, some of which may be unreliable (e.g., open repositories like HuggingFace Lhoest et al. (2021)). As a result, models can unknowingly incorporate poisoned data, leading to backdoor vulnerability Zhang et al. (2021); Xu et al. (2021): a backdoored model, trained with poisoned data, behaves normally on clean inputs but misbehaves when exposed to the trigger Bai et al. (2025).
A critical yet underexplored problem arises when such backdoors are discovered after model deployment. Existing defenses, such as data filtering and retraining (He et al., 2024), pruning followed by fine-tuning Liu et al. (2018); Zhao et al. (2024b), or unlearning backdoors based on clean Li et al. (2023) or poisoned data Min et al. (2024); Li et al. (2021b); Chen et al. (2022), typically incur substantial computational costs Wu et al. (2022)–especially for large-scale models or proprietary pipelines. This raises a practical and timely question: Can we effectively purify a trained backdoored model without retraining?
Recent work (Arora et al., 2024) has proposed model merging as a cost-efficient and retraining-free defense method for model purification: leveraging the abundance of open-source proxy models online [^1] For reference, on Hugging Face, there are 74 LLaMA-7B and 262 BERT-base models trained on SST-2 alone, and the pool of proxy models is even larger when considering all datasets within the same task domain., this approach merges several proxy models–backdoored or not–with the victim model. While promising, this approach relies on ad hoc merging of multiple models with no principled selection, such as utility and defense performance. Its effectiveness is thus unpredictable and sensitive to the number of models involved. Our later experiments in Section˜5.2 demonstrate that when only a single proxy model is available, such defenses become ineffective. More seriously, In practice, collecting more homogeneous proxy models also requires more effort and often harm utility more Zhou et al. (2025a); Wang et al. (2025b), even though more merged parameters can help neutralize backdoors. We therefore target a more challenging scenario–designing guided merging using only one proxy model.
In this work, we propose a simple yet effective alternative: guided merging with a single proxy model. Rather than blindly merging models, we explicitly guide the merging process using a signal that approximates the trade-off between utility and backdoor risk. Intuitively, if retaining a particular module from the victim model contributes significantly more to the attack performance than to task utility, it is likely critical for encoding backdoor features. We thus progressively identify such modules and replace them with their counterparts from the proxy model, as long as the resulting utility degradation remains limited relative to backdoor mitigation gain. We refer to this method as Guided Module-Substitution (GMS). While training-free and more efficient than prior model-merging-based defenses, our method GMS not only retains the robustness of model-merging defenses to proxy model trustworthiness, but is also tolerant of imperfect data knowledge, stable across hyperparameters, and exhibits strong transferability across different attacks. Further discussion of these desirable properties is in Section˜5.3.
The contributions of this work are as follows.
  1. 1.
    We propose GMS, the first backdoor purification method based on guided merging of a victim model with a single proxy model. This design enables purification without retraining while remaining efficient and practical.
  2. 2.
    Experiments on standard NLP models and LLMs (Section˜5.2) show that GMS consistently surpasses competitive defense baselines, with especially strong performance against the more challenging attacks like LWS and HiddenKiller.
  3. 3.
    We further validate four desirable properties of GMS in Section˜5.3 and Sections˜B.4, B.5, B.6, B.7, B.8, B.9, B.10, B.11 and B.12, namely (1) robustness to proxy model choices, (2) applicability to varied data conditions, (3) stability across hyperparameters, and (4) transferability across attacks.
2 Related works
Backdoor attack.
Backdoor attacks aim to manipulate models to behave normally on benign inputs while exhibiting attacker-controlled behavior when specific triggers are present. They can be categorized into two series Wu et al. (2022): (1) Data-poisoning attacks, exemplified by BadNets Gu et al. (2017), involve tampering with a subset of training data by adding triggers and altering labels, making them practical in real-world scenarios with minimal attacker assumptions Turner et al. (2019); Carlini and Terzis (2022); Goldblum et al. (2023); Li et al. (2021c). Initially explored in image classification, backdoor attacks have since raised significant security concerns in NLP Dai et al. (2019); Qi et al. (2021c, b) with triggers range from misspelled words Chen et al. (2021), rare words Yang et al. (2021a); Kurita et al. (2020) and syntactic structures Qi et al. (2021b) to text styles Pan et al. (2022), posing serious challenges for detection. (2) Training-control attacks, on the other hand, assume complete control over the training process and data Nguyen and Tran (2021); Kurita et al. (2020), with advances such as layer-wise weight poisoning ensuring backdoor persistence even after fine-tuning Li et al. (2021a). In this work, we focus on data-poisoning backdoor attacks in NLP due to their widespread applicability and practical implications.
Backdoor defense.
Post-training backdoor defense approaches can be broadly categorized as: (1) Backdoor sample detection followed by training, which first filters poisoned samples from the dataset and then retrains or fine-tunes the model. These defenses typically rely on the assumption that poisoned samples exhibit distinct characteristics compared to clean (benign) ones, enabling their detection Qi et al. (2021a); Yang et al. (2021b); Tran et al. (2018); He et al. (2023), or attempt to reverse-engineer backdoor triggers to neutralize their impact Wang et al. (2019, 2022, 2023); Xu et al. (2024). Afterward, such strategies usually involve unlearning the identified backdoor samples or retraining/fine-tuning on the filtered dataset. However, their effectiveness is often unsatisfactory Sun et al. (2025); Wang et al. (2023); Qi et al. (2021a), primarily because accurately identifying all poisoned samples is (increasing) difficult Wu et al. (2022); Qi et al. (2021a); Yang et al. (2021b); Tran et al. (2018); He et al. (2023), leaving residual backdoor behavior even after retraining or unlearning. Our method, in this regard, compensates for these shortcomings by being more robust to data filtering quality, and thus provides a more reliable alternative to unlearning- or retraining-based defenses. (2) Backdoor model purification: These methods aim to eliminate backdoor features directly from a well-trained model and are generally regarded as achieving state-of-the-art defense performance Zhu et al. (2023); Zhao et al. (2024b). Common approaches involve merging parameters with other proxy models Arora et al. (2024); Chen et al. (2024a), or pruning and fine-tuning using a limited amount of clean data Wu and Wang (2021); Min et al. (2023); Liu et al. (2018); Zhao et al. (2024b); Zhu et al. (2023), which may sometimes be assisted by clean proxy models as well Liu et al. (2018); Zhang et al. (2023). Our work advances the second line of research by minimizing the number of auxiliary models to one and by eliminating the need for additional training and strictly clean datasets.
Model merge.
Model merging methods have emerged as an efficient way to build powerful models by combining existing trained domain-specialized models without requiring extensive retraining Wortsman et al. (2022); Ilharco et al. (2023); Cheng et al. (2025). Most advances focus on developing utility-preserving multi-task models (e.g., merging models trained on different tasks like coding and mathematics He et al. (2025)), or stronger domain-specific models by combining within-domain models Wortsman et al. (2022). Surprisingly, recent work shows that naive weight averaging across multiple models can mitigate backdoors Arora et al. (2024). Yet such blind merging is inefficient–requiring unpredictable trial-and-error in model collection, since one cannot know in advance how many models are needed for purification–and often harmful, as merging many models can yield utility drops from parameter interference and redundancy Zhou et al. (2025a); Wang et al. (2025b). Thus, our work address these issues by proposing effective and efficient guided merging with a single proxy model. Note that while there are some works studying how model merging systems themselves can be exploited for backdooring, leveraging the fact that merging methods aim to preserve (task) parameters from all models Wang et al. (2025a); Zhang et al. (2024); Yuan et al. (2025), our objective is fundamentally different: we use merging as a tool to purify a single target model, where only the target model’s clean task is relevant and other tasks can all be discarded. Since our main goal is to demonstrate guided merging can act as a defense–and obtaining a single homogeneous proxy is practical–we focus on the standard homogeneous setting. As heterogeneous merging itself remains an open challenge Xu et al. (2025); Du et al. (2025), exploring its use for defense is left to future work.
Safety Localization.
Recent studies suggest that safety-critical behaviors in LLMs are often localized to specific layers or components. For example, jailbreak defenses identify safety layers that are critical for aligning harmful queries Zhao et al. (2024a); Ouyang et al. (2025); Zhou et al. (2024); a small set of safety layers Li et al. (2025) or safety modules (e.g., attention heads) Zhou et al. (2025b) are crucial for distinguishing malicious from benign inputs; and modifying specific safety-related parameters can further enhance alignment Wei et al. (2024). Similarly, backdoor research shows that certain layers Jebreel et al. (2023) and even neurons or heads Zhao et al. (2024b) disproportionately influence attack success, enabling defenses through targeted editing or masking. These findings motivate our focus on module- and layer-level guided merging, which balances efficiency with effectiveness.
3 Preliminary
Modern neural network (NN) architectures can be viewed as layer-wise compositions of functional blocks, where each block serves as a reusable unit performing a well-defined computational operation. Hereby, we define two aspects to systematically describe generic NN architectures:
  1. 1.
    Layer Set ($\mathcal{L}$) is the set of block indices that captures the depth of the model.
  2. 2.
    Module Set ($\mathcal{M}$) is the set of representative functional modules within each block.
Transformer blocks.
In this paper, we use Transformer-based architecture Vaswani et al. (2017) as the testbed, given its widespread attention and implementation in NLP.
For each Transformer block, the input will be processed through Attention and Feed-Forward Network (FFN) modules. An attention module contains four components: $W_{Q}$, $W_{K}$, $W_{V}$, and $W_{O}$, projecting the input into query, key, value, and the final output representation, respectively. The following FFN contains two components: $W_{F}$ for the forward layer and $W_{P}$ for the projection layer, connected by a non-linear activation function.
In total, each Transformer block contains six key functional modules with parameters $W_{Q},W_{K},W_{V},W_{O},W_{F},$ and $W_{P}$. For simplicity, we adopt notations $Q,K,V,O,F,$ and $P$ to refer to these weight matrices throughout the paper. Accordingly, we define the module set for Transformer models $\mathcal{M}=\{Q,K,V,O,F,P\}$ and layer set $\mathcal{L}=\{1,2,\ldots,L\}$.
Figure 1: The pipeline of our method. Step 1: Extract two small proxy datasets used for computing score in Equation˜1. Step 2: Iteratively update a substitution matrix $S$ to greedily maximize the trade-off score between backdoor removal and utility preservation. Step 3: Return purified model $M_{pure}$ corresponding to the best substitution matrix $S_{best}$. For more details, refer to Algorithm˜1.
4 Methodology
In Section˜4.1, we introduce the problem setting, and we present our methods in Section˜4.2.
4.1 Problem setting
Defense Setting.
The defender (e.g., a model vendor) has trained a potentially backdoored victim model $M_{v}$ on an unauthenticated dataset $\mathcal{D}$, which may contain poisoned samples. Following the standard setting Li et al. (2021b), we assume the defender has full control over the training process but lacks prior knowledge of the backdoor patterns, their proportion, or their distribution within $\mathcal{D}$. If an isolation method is employed, it may only identify a subset of the poisoned examples, as perfect detection is not guaranteed. The defender has white-box access to the trained victim model $M_{v}$, as well as access to a proxy model $M_{p}$ trained on similar downstream tasks.
4.2 Our method
We aim to combine a victim model with a proxy model. Unlike traditional model merging Arora et al. (2024), which prioritizes utility across all models’ tasks, our method targets backdoor removal and requires preserving only the clean task performance of the target model. This enables a more targeted substitution strategy: directly replacing backdoor-encoded modules in the victim model $M_{v}$ with counterparts from the proxy model $M_{p}$. Specifically, we operate at the module- and layer-level, inspired by recent findings that malicious behaviors in LLMs are often localized to specific safety layers and safety modules (discussed in Section˜2). While not the only possible granularity, this level strikes a balance between efficiency and effectiveness, and offers greater robustness and transferability than neuron-level edits, as discussed in Section˜4.2.3 and Section˜5.3.
Our main algorithm, Guided Module-Substitution (GMS), is detailed in Section˜4.2.3. Section˜4.2.1 defines the objective optimized by GMS, and Section˜4.2.2 provides an example of constructing proxy datasets.
4.2.1 Objective
The defender’s objective is to identify a purified model $M$ within the model space $\mathbb{M}$, defined by all possible substitution strategies over the layer set $\mathcal{L}$ and module set $\mathcal{M}$, that effectively eliminates the backdoor while preserving task utility:
$\operatorname*{argmax}_{M}\,(1-\alpha)\cdot\Delta_{\text{asr}}(M)+\alpha\cdot(1-\Delta_{\text{acc}}(M))$ (1)
where $\Delta_{\text{asr}}(M)$is a score that reflects the extent of backdoor removal and $\Delta_{\text{acc}}(M)$ assesses the reduction in task utility. Specifically, if given access to clean and poisoned samples, we can evaluate the attack success rate (ASR) on poisoned data and the clean accuracy (ACC) on benign inputs. The improvement in backdoor defense and degradation in utility for a given model $M$ relative to the original victim model $M_{v}$ are defined as:
$\Delta_{\text{asr}}(M)$ $=s_{asr}(M_{v})-s_{asr}(M),$ (2)
$\Delta_{\text{acc}}(M)$ $=s_{acc}(M_{v})-s_{acc}(M).$ (3)
The hyperparameter $\alpha\in[0,1]$ controls the trade-off between backdoor removal and task utility preservation. A larger $\alpha$ favors utility, while a smaller $\alpha$ prioritizes backdoor defense.
4.2.2 Proxy datasets $\mathcal{D}_{poison}$ and $\mathcal{D}_{clean}$
To approximate Equations˜2 and 3, we extract two proxy datasets: a proxy clean set $\mathcal{D}_{\text{clean}}$ and a proxy poison set $\mathcal{D}_{\text{poison}}$, without assuming any prior knowledge of the backdoor. We only follow the standard setting to assume access to the (potentially poisoned) training set $\mathcal{D}$, or to a significantly smaller subset (e.g., 1%) of it (e.g., in user-reporting scenarios).
In literature, various methods can be used to construct proxy datasets under such settings, such as sample diagnosis Gao et al. (2019); Huang et al. (2022); Guo et al. (2023) and trigger inversion Wang et al. (2019, 2022, 2023); Xu et al. (2024). For simplicity, we adopt the following heuristics in this paper: we construct $\mathcal{D}_{\text{clean}}$ via random sampling, assuming that poisoned samples do not dominate the dataset; and we construct $\mathcal{D}_{\text{poison}}$ using a naive confidence-based heuristic, based on the observation that poisoned examples often yield higher output confidence scores on backdoored models Li et al. (2021b); Swayamdipta et al. (2020).
4.2.3 Greedy search for purified model $M_{pure}$
Using the extracted proxy datasets $\mathcal{D}_{\text{poison}}$ and $\mathcal{D}_{\text{clean}}$, we compute (potentially inaccurate) estimates of Equations˜2 and 3222An imperfect $\mathcal{D}_{\text{poison}}$ will lead to an underestimated $\Delta_{\text{asr}}(M)$, while an impure $\mathcal{D}_{\text{clean}}$ results in an overestimated $\Delta_{\text{acc}}(M)$. However, what matters is only the relative scaling between these two after weighting by $\alpha$ (Appendix C).and plug into the objective Equation˜1 to get the scoring function that guides the substitution process.
While our goal is to purify $M_{v}$ by replacing its modules across $L$ layers, an exhaustive search is computationally impractical due to its exponential complexity (e.g., $2^{6\times 12}$ for a BERT-base model (Devlin et al., 2019)). Moreover, considering individual parameters as functional units for editing often falls short in breaking backdoor connections (thereby leading prior purification works to heavily rely on clean fine-tuning to restore clean features Liu et al. (2018)), we propose a greedy algorithm that transforms the problem into a feasible search–balancing granularity and scalability by iteratively identifying and replacing the most critical module, maximizing the objective score to derive the optimal purified model $M_{pure}$.
Algorithm 1 Guided-Module-Substitution (GMS)
1:Input: Victim model $M_{v}$, proxy model $M_{p}$, module set $\mathcal{M}$, layer set $\mathcal{L}$, proxy datasets $\mathcal{D}_{clean}$ and $\mathcal{D}_{poison}$
2:Output: Purified model $M_{pure}$
3: Initialize $S\leftarrow\mathbf{1}_{|\mathcal{M}|\times|\mathcal{L}|}$, $c_{best}\leftarrow$-$\infty$, $S_{best}\leftarrow\varnothing$
[…]
5:# Iteratively Update Selected Modules and Layers:
6:while $|\mathcal{M}|>1$ or $|\mathcal{L}|>1$ do
7:  $c^{*}\leftarrow$-$\infty$, $S^{*}\leftarrow\varnothing$
8:  for each $m\in\mathcal{M}$ or $l\in\mathcal{L}$ do
9:   $S^{\prime}\leftarrow S$ with $S^{\prime}_{:,m}=\mathbf{0}_{L\times 1}$ or $S^{\prime}_{l,:}=\mathbf{0}_{1\times|\mathcal{M}|}$
10:   $M_{pure}^{\prime}\leftarrow$ Replace($M_{v}$, $M_{p}$, $S^{\prime}$)
11:   $c\leftarrow$ ComputeScore($M_{pure},s_{acc},s_{asr}$)
12:   if $c>c^{*}$ then
13:    Update $c^{*}\leftarrow c,S^{*}\leftarrow S^{\prime}$
14:   end if
15:  end for
16:  $S\leftarrow S^{*}$ and update $\mathcal{M},\mathcal{L}$ accordingly
17:  if $c^{*}>c_{best}$ then
18:   Update $c_{best}\leftarrow c^{*}$, $S_{best}\leftarrow S^{*}$
19:  end if
20:  if $c_{best}$ not been updated for $T$ iterations then
21:   break
22:  end if
23:end while
24:return $M_{pure}=\text{Replace}(M_{v},M_{p},S_{best})$
Purifying $M_{v}$ by module substitution.
Given a victim model $M_{v}$ (likely to has high ACC and high ASR), we next illustrate how to identify the specific modules and layers that, when substituted, allow the model to maintain strong performance on the clean task (i.e., minimizing $\Delta_{\text{acc}}$) while eliminating the backdoor features (i.e., maximizing $\Delta_{\text{asr}}$).
Our algorithm, as demonstrated in Figure˜1 and Algorithm˜1, tracks the parameters selected for substitution using a module set ($\mathcal{M}$) and a layer set ($\mathcal{L}$), e.g., if $\mathcal{M}=\{O,F,P\}$ and $\mathcal{L}=\{7,8,9\}$, it indicates that the $O$, $F$, and $P$ module parameters in the 7th, 8th, and 9th layers of the victim model $M_{v}$ should be replaced with the corresponding parameters from proxy model $M_{p}$. For simplicity, we introduce a substitution matrix $S\in\{0,1\}^{|\mathcal{M}|\times|\mathcal{L}|}$ to specify which modules in which layer in the victim model $M_{v}$ are selected for substitution. Note that while the shape of $S$ is determined by the initial sizes of sets $\mathcal{M}$ and $\mathcal{L}$, these two sets are iteratively updated throughout the algorithm. Each row of $S$ corresponds to a layer in the layer set $\mathcal{L}$, and each column corresponds to a module in the module set $\mathcal{M}$. The value of $S[l,m]=1$ indicates that the module $m\in\mathcal{M}$ in layer $l\in\mathcal{L}$ in the victim model $M_{v}$ is selected for substitution, and $S[l,m]=0$ otherwise. Given $S$, we can succinctly represent the parameter substitution as $S\odot M_{p}+(1-S)\odot M_{v}$, where $\odot$ denotes the element-wise product.
Since the primary goal is backdoor removal, we start by selecting all modules in $M_{v}$ to be replaced with those of $M_{p}$, i.e., $\mathcal{M}=\{K,Q,V,O,F,P\}$, $\mathcal{L}=\{1,\ldots,L\}$, $S=\mathbf{1}_{|\mathcal{M}|\times|\mathcal{L}|}$. At $t$-th iteration, the algorithm considers two types of updates to $S^{t}$: deselecting a module from the current module set $\mathcal{M}$ (affecting all selected layers) or deselecting a layer from the layer set $\mathcal{L}$ (affecting all selected modules). The update that maximizes the score in Equation˜1 is applied to $S^{t+1}$. The process continues until one of two conditions is met: (1) further iterations will not yield a better strategy (controlled by stopping patience $T$); or (2) no additional layers or modules can be removed, i.e., $|\mathcal{M}|=|\mathcal{L}|=1$. Finally, the strategy $S$ that yields a purified model $M_{pure}$ with the highest score is returned. The complexity of our algorithm is a practical quadratic $O(|\mathcal{M}|^{2}+|\mathcal{L}|^{2})$.
5 Experiments
5.1 Experimental settings
Datasets. We evaluate our method on four datasets: SST-2 (Socher et al., 2013), OLID (Zampieri et al., 2019), MNLI (Williams et al., 2018), and AGNEWS (Zhang et al., 2015). These datasets cover text classification and natural language inference (NLI) tasks, including binary and multi-class classification scenarios, and are widely used for evaluating text backdoor attacks and defenses (Qi et al., 2021a; He et al., 2023; Gupta and Krishna, 2023; Arora et al., 2024). We adapt the open-source datasets provided by HuggingFace (Lhoest et al., 2021). Table 1 are the statistics of these datasets.
Backdoor Attacks.
We study defenses against four prominent types of data-poisoning backdoor attacks: (1) BadNets (Kurita et al., 2020), (2) InsertSent (Dai et al., 2019), (3) Learnable Word Substitution (LWS(Qi et al., 2021c), and (4) HiddenKiller (Qi et al., 2021b). The first two correspond to insertion-based methods, while the latter two involve synonym substitution and syntactic paraphrasing approaches, respectively. A more recent attack, BITE Yan et al. (2023), is discussed separately in Section˜B.3 due to its atypical behavior on benign models.
Follow Arora et al. (2024), we use rare words {"cf", "mn", "bb", "tq", "mb" } as triggers for BadNets and phrases {"I watched this movie", "no cross, no crown" } for InsertSent. The poison target labels for each dataset are listed in Table 1. While our main experiments focus on 20% poison rate (in line with the literature (Dai et al., 2019; Qi et al., 2021b; Arora et al., 2024)), i.e., 20% of training samples are poisoned, we also evaluate lower poison rates of 10%, 5% in Section˜B.4.
Dataset Classes Train Test Target Class
Clean Poison
SST-2 2 67,349 872 444 Negative (0)
OLID 2 13,240 860 240 Not offensive (1)
MNLI 3 100,000 400 285 Neutral (1)
AGNews 4 120,000 7,600 5,700 Sports (1)
Table 1: The statistics of our evaluated datasets.
BadNet InsertSent LWS HiddenKiller
[…]
Benign 4.1 95.9 2.2 95.9 12.8 95.9 16.5 95.9
Victim 100.0 96.0 100.0 96.3 98.0 95.4 96.5 95.7
Proxy Model (IMDB) 7.4 89.1 4.3 89.1 10.8 89.1 13.7 89.1
ONION 56.8 92.9 99.9 93.3 85.7 91.9 92.9 92.9
Z-Def. 4.6 96.1 1.8 95.6 97.3 95.3 35.7 95.4
PURE 0.0 50.9 0.0 50.9 0 50.9 0.0 50.9
ABL 75.0 49.2 51.7 50.8 32.3 50.3 92.9 47.4
TIES 99.9 95.7 100.0 95.8 93.5 95.2 88.8 95.3
DARE w/ TIES 99.3 96.0 100.0 96.2 96.4 95.7 92.7 95.8
WAG 84.4 94.8 60.1 95.2 58.8 94.8 56.2 92.5
SST-2 Ours (GMS) 4.5 91.6 1.9 92.5 9.7 91.7 10.4 91.2
Benign 1.9 95.4 0.5 95.4 0.5 95.4 1.1 95.4
Victim 99.9 95.1 99.6 95.3 99.6 94.5 100.0 95.1
Proxy Model (BBCNews) 1.5 70.2 1.7 70.2 1.8 70.2 3.4 70.2
ONION 59.4 94.8 97.8 95.1 84.8 94.5 99.6 94.7
Z-Def. 1.6 95.3 0.4 95.3 97.9 96.1 100.0 95.0
PURE 2.8 86.3 3.0 85.4 6.5 85.0 9.4 84.6
ABL 50.2 55.0 48.8 54.8 100.0 55.0 90.2 54.9
TIES 99.9 94.6 99.6 94.4 97.7 95.8 100.0 94.4
DARE w/ TIES 99.9 95.2 99.6 95.4 97.8 96.5 100.0 95.2
WAG 92.7 94.1 97.8 94.4 78.0 93.9 90.9 94.3
AGNews Ours (GMS) 2.5 91.0 2.4 92.6 3.2 91.7 6.5 90.4
Table 2: (Partial) Performance of our method on two datasets compared to baselines under various backdoor attacks on the RoBERTa-large model, with each value averaged over three seeds. We highlight the top-2 lowest ASR results in blue cells, and the highest CACC results in bold. Results for other datasets are in Table˜8.
Defense Baselines.
We choose seven renowned defenses as main baselines, including two data-wise detection methods: (1) ONION (Qi et al., 2021a) and (2) Z-Def. (He et al., 2023), and four model-wise purification methods: (3) PURE (Zhao et al., 2024b), (4) ABL (Li et al., 2021b), (5) TIES (Yadav et al., 2023), (6) DARE (Yu et al., 2024), and (7) WAG (Arora et al., 2024).
Among these, ONION and Z-Def. detect and remove triggers from datasets. ONION identifies outlier words (potential triggers) using language models, e.g., GPT-2 (Radford et al., 2019), while Z-Def. detects spurious correlations between tokens and labels. In contrast, (2)-(7) are model purification methods. PURE purifies the victim model using attention head pruning and normalization techniques, and ABL introduces a robust anti-backdoor training framework that first isolates, then unlearns backdoor associations. TIES, DARE and WAG are mode-merging baselines. We also report three additional baselines in Section˜B.2: two model purification methods, Fine-mixing (Zhang et al., 2022) and Fine-purifying (Zhang et al., 2023), and one data filtering method, SEEP (He et al., 2024), followed by retraining.
Evaluation Metrics.
Following previous literature (Qi et al., 2021a), we adopt Clean Accuracy (CACC) and Attack Success Rate (ASR) to measure utility and defense performance respectively. CACC is evaluated on a clean test set, where both the samples and labels are ground truth. In contrast, ASR is evaluated on a poisoned test set, where triggers are implanted into each sample, and the labels are flipped to the target class.
Implementation details.
Following Arora et al. (2024), we compare all methods on RoBERTa-large (Liu et al., 2019), BERT-base-uncased (Devlin et al., 2019), as well as LLMs including Llama 2 7B (Touvron et al., 2023), Mistral 7B (Jiang et al., 2023), and Qwen 2.5 7B (Team, 2024), with low-rank adaptation LoRA (Hu et al., 2022). All victim models are fine-tuned on poisoned datasets using the Adam optimizer with no weight decay (Kingma and Ba, 2015) and a learning rate of $2\times 10^{-5}$. All defense baselines are implemented based on their open-source repositories (see Section˜A.3). The high-confidence proxy datasets ($\mathcal{D}_{clean}$ and $\mathcal{D}_{poison}$) are extracted following Section˜4.2.2. To train proxy models for selection, we use IMDB (Zhang et al., 2015) for the SST-2 victim model, Twitter Abusive (Founta et al., 2018) for the OLID, SNLI (Young et al., 2014) for MNLI, and BBCNews (Greene and Cunningham, 2006) for AGNews.
Our method uses hyperparameter $\alpha$ in Equation˜1 to control the trade-off between backdoor defense and task utility. By default, $\alpha$ is set to 0.4 but can be adjusted based on priorities: smaller values (e.g., 0.1) for lower ASR and larger values (e.g., 1.0) for higher utility. The stopping patience $T=5$ is default in Algorithm˜1. In Section˜B.5, we justify these choices and show that our method is NOT particularly sensitive to hyperparameters.
5.2 Main results
Table˜2 presents the performance of our method compared to all baselines under benchmark backdoor attacks on the RoBERTa-large model, with scores averaged over three runs using different seeds. Results on other datasets and architectures are provided in Section˜B.1. Across all datasets and backdoor attacks, our method consistently ranks among the top 2, regarding backdoor removal performance, with minor harm to clean accuracy. In particular, under the two particularly challenging backdoor tasks, LWS and HiddenKiller, GMS demonstrates significant improvements (i.e., at least 25%) over all baselines–for example, for SST-2, our method reduces the ASR on LWS to 9.7%, compared to 58.8% for the following baseline Z-Def. While Z-Def achieves competitive performance with GMS in defending against BadNet and InsertSent, Z-Def is much less effective against LWS on both datasets. We attribute this to Z-Def.’s reliance on lexical and syntactic features to detect outliers in the poisoned dataset, whereas LWS attacks subtly replace words with synonyms, effectively bypassing outlier detection. As for PURE, another competitive recent approach, we found its head-pruning step is highly unstable across both tasks and architectures. For example, it performs well on the BERT-base model in Table˜9 (following the settings reported in PURE (Zhao et al., 2024b)), but performs poorly on most tasks with RoBERTa-large. Unlearning-based methods (ABL Li et al. (2021b)) exhibit similar issues–often degrading CACC more than reducing ASR. Such instability has also been noted in prior works Liu et al. (2018); Wu et al. (2022), which shows that these methods are sensitive to both hyperparameter choices and the quality of the proxy data. Additionally, we observed that all model-merging baselines suffer in defense performance when merging with a single proxy model, aligning with the results in Arora et al. (2024), and tend to prioritize preserving accuracy over removing backdoors. We discuss the potential reasons for this and the instability of PURE in Section˜B.1. We also compare our approach with two earlier model purification baselines, Fine-mixing Zhang et al. (2022) and Fine-purifying Zhang et al. (2023) in Section˜B.2.
Performance on LLMs.
Given the growing popularity of LLMs, we evaluate the performance of our method on Llama-2-7B, Mistral-7B and Qwen-2.5-7B for the SST-2 dataset (Table˜4). Remarkably, in all cases, GMS effectively removes backdoors.
[…]
Hidden Killer 95.4 4.5
BadNet (Diff. Trigger) 87.8 6.9
BadNet (Same Trigger) 95.8 100.0
Victim Model (BadNet) 95.6 100.0
Table 3: Proxy models with implanted backdoors can still effectively purify victim models if the backdoor is not identical in both attack strategy and trigger.
5.3 Ablation studies
We next examine the key property of GMS–robustness across diverse proxy datasets. For completeness, we defer additional results to the appendix: stability under different hyperparameter settings (Section˜B.5), transferability across attacks (Section˜B.12), preservation of clean victim model utility (Section˜B.6), and resilience to re-tuning attacks (Section˜B.10).
BadNet InsertSent LWS HiddenKiller
[…]
Benign 3.6 96.7 3.5 96.7 14.9 96.7 15.2 96.7
Victim 100.0 97.0 100.0 97.1 98.7 96.1 95.9 96.7
Llama-2 GMS 3.5 91.7 3.6 93.4 12.2 91.5 7.6 91.1
Benign 3.8 97.1 4.5 97.1 7.7 97.1 12.3 97.1
Victim 100.0 95.8 100.0 96.5 98.9 95.5 95.9 96.4
Mistral GMS 4.4 92.1 7.1 92.4 3.8 91.9 5.3 91.9
Benign 5.0 97.0 2.7 97.0 15.1 97.0 15.7 97.0
Victim 100.0 96.3 100.0 92.9 98.5 98.5 96.0 96.2
Qwen-2.5 GMS 6.3 94.5 7.5 94.4 16.4 91.7 15.0 93.0
Table 4: Performance of our method on Llama-2-7b, Mistral-7b and Qwen-2.5-7b on SST-2 dataset, averaged over three seeds.
BadNet InsertSent LWS HiddenKiller
Victim Model Proxy Model ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$
IMDB 5.4 92.8 1.6 94.0 12.6 93.8 11.0 93.0
YELP 4.3 96.8 8.1 94.8 23.2 92.3 19.1 93.8
SST-2 AMAZON 3.8 92.9 1.1 95.8 20.1 92.6 13.5 93.8
Table 5: Proxy models trained on different datasets can all effectively purify victim models (trained on the SST-2).
Proxy-model robustness:
GMS retains the robustness property of model-merging defenses: it is resilient to both the specific choice of the proxy model and the benign or malicious nature of the proxy model; i.e., as long as the backdoors in the two models are not identical, the backdoor can be effectively removed. We examine the sensitivity of GMS to the selection of proxy models in two scenarios. Scenario 1: Different proxies. We evaluated the performance of using different proxy models trained on homologous datasets, such as IMDB, Yelp, and Amazon (Zhang et al., 2015), to purify a victim model trained on SST-2. As shown in Table˜5, our method is largely insensitive to the choice of proxy model across all backdoor attacks. Using any proxy model, GMS effectively defends against all tested attacks. Scenario 2: Backdoored proxies. We tested cases where the victim model and the proxy model were trained on the same dataset (SST2) with different backdoor attacks (BadNet and Hidden Killer) or on different datasets (SST2 and IMDB) with the same attacks (BadNet). From Table˜3 and Table˜17, we observed that as long as the backdoor in proxy model is not identical to that in victim model (i.e., the same backdoor strategy with an identical backdoor trigger), GMS consistently mitigates the backdoor. We attribute this to the combination of module-level substitution and disruptive nature of model merging, which breaks the backdoor pathway such that a non-identical proxy cannot restore it.
Sub-proxy datasets GMS
Proxy Clean Proxy Poison CACC $\uparrow$ ASR $\downarrow$
Random Sample Outlier Detection 93.6 3.2
Random Sample Oracle 94.9 4.1
Oracle Outlier Detection 92.7 4.3
Oracle Oracle 95.6 4.3
Benign SST-2 95.6 4.1
Table 6: Performance comparison of different sub-proxy dataset configurations on GMS.
Proxy-data robustness:
Prior pruning- and unlearning-based methods strictly rely on access to clean data or known backdoor triggers Min et al. (2024); Li et al. (2021b); Chen et al. (2022); Li et al. (2023), typically extracted from the training subset. In contrast, GMS operates under more practical settings, e.g., post-processing after suboptimal data detections, where the data received is a mixture of clean and poisoned samples. Our previous experiments used $\mathcal{D}_{clean}$ from random sampling (containing 20% poisoned data) and $\mathcal{D}_{poison}$ from heuristic outlier detection. Table˜7 shows that even when 70% of the “clean” proxy dataset is poisoned, GMS is still effective in purification. In Appendix˜C, we mathematically justify this and discuss the analytical constraints on proxy datasets. These results demonstrate that our method is highly robust to proxy-data construction. Nevertheless, Table˜6 shows that more accurate proxy datasets–such as an oracle clean or label-flipped poison dataset–can further enhance performance, enabling precise module identification and full utility preservation (e.g., CACC of 95.6%). However, even without optimal proxies, our method consistently removes backdoors, achieving benign-level ASR (4.1%) with comparable CACC.
SST-2 AGNews
Ratio $\rho$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$
0.00 93.58 3.38 91.76 2.42
0.30 91.51 3.15 91.59 3.61
0.60 89.11 5.18 88.67 1.42
0.70 89.11 5.18 88.67 1.42
0.80 89.11 5.18 94.59 99.75
0.90 89.11 5.18 94.83 99.82
Table 7: Purification performance as the ratio $\rho$ of poisoned data in extracted $\mathcal{D}_{clean}$ increases.
6 Conclusion
We propose Guided-Module-Substitution (GMS), a model purification defense against data-poisoning attacks in NLP based on guided merging of the victim model with a single (existing) proxy model. Extensive experiments on both standard NLP models and LLMs validate its effectiveness, particularly against challenging attacks like LWS and HiddenKiller. Beyond strong empirical results, we validated desirable properties of GMS: robustness to proxy selection, tolerance to imperfect data, stability across hyperparameters, and transferability across attacks. These findings suggest that guided single-proxy merging offers a practical alternative to retraining-based defenses.
Limitations
Our method advances model-merging-based defenses by guiding the merging process with a trade-off signal between task utility and backdoor risk, reducing the requirement from a number of proxy models to just one. One potential limitation is that, like prior proxy-model-based defenses, we focus on the homogeneous setting and do not explore heterogeneous proxies. We make this choice for two reasons. First, in the defense scenario, merging serves solely as a tool for purification rather than an end in itself, so defender can control which proxy model to use. In practice, obtaining a homogeneous proxy is not difficult, making the heterogeneous case less relevant for our scenarios. Second, our goal is to introduce and validate the concept of guided merging for backdoor defense, rather than to develop state-of-the-art merging techniques. Since knowledge transfer across heterogeneous models itself remains an open challenge (Xu et al., 2025; Du et al., 2025), we leave such extensions to future work. Moreover, while our method tolerates the use of a malicious proxy model, as long as its backdoor is not identical to the victim’s (i.e., the same attack method with the exact same trigger, which is rare in practice), as shown in Table˜17, further relaxing this, such as enabling defenses against identically backdoored proxies or developing proxy-free approaches, could lead to even more robust and powerful defense strategies.
Ethical statement
Our method introduces a retraining-free purification approach for backdoored models, contributing to mitigating security risks in NLP systems from backdoor attacks. Additionally, our defense approach reduces the need for reannotating datasets to ensure they are purely clean, minimizing the potential harm of exposing annotators to harmfully poisoned contents. While we do not anticipate any direct negative consequences from this work, we hope it inspires further advancements in the development of robust, retraining-free defense methods for more realistic scenarios in future research.
Acknowledgements
We acknowledge the support from 2024 FSE Strategic Startup and the credits awarded from Google Cloud Platform.
Appendix A Experiment details
A.1 Experiment setup
We adopt the same training configuration as WAG (Arora et al., 2024) to ensure a fair comparison. This includes fine-tuning RoBERTa-large, BERT-base-uncased, as well as LLMs including Llama 2 7B (Touvron et al., 2023), Mistral 7B (Jiang et al., 2023), and Qwen 2.5 7B (Team, 2024) on poisoned datasets, using the Adam optimizer with no weight decay (Kingma and Ba, 2015) and a learning rate of $2\times 10^{-5}$. For encoder models, the batch size, maximum sequence length, and epoch are set to 32, 128, and 3, respectively. For decoder LLMs, due to computational limits, we adapt low-rank adaptation (LoRA) (Hu et al., 2022) for all the linear modules within the Transformer layers and apply our method to substitute those LoRA modules after training. We train these LLMs for 2 epochs with a batch size of 8 and a maximum sequence length of 128.
To validate our defense, we first adapt the aforementioned backdoor methods and datasets to obtain victim models. Specifically, to train backdoored models, we construct poisoned datasets by injecting triggers into the training split, while the validation split is used to construct a clean and a poison test set respectively. For MNLI, we use a random subset of 100,000 samples to reduce overhead. Next, we train proxy models on homologous datasets to prepare for future merging. We then construct high-confidence proxy-clean ($\mathcal{D}_{clean}$) and proxy-poison ($\mathcal{D}_{poison}$) sets from the training data (using the methods illustrated in Section˜4.2.2) to guide our parameter substitution strategy search.
For a model trained on unverified data, access to fully clean data is often not guaranteed. However, the corresponding task is typically known to the defender, providing an opportunity to identify homologous models trained on overlapping tasks that can be used for merging. For instance, previous works, WAG and PURE, utilized the IMDB (Maas et al., 2011) dataset to purify models trained on poisoned SST-2 datasets, as both share the sentiment classification task domain.
In this work, we select proxy models that share a similar domain with the poisoned tasks. Specifically, we use Twitter Abusive (Founta et al., 2018) for OLID, SNLI [^3] Creative Commons Attribution-ShareAlike 4.0 International License (Young et al., 2014) for MNLI, and BBCNews (Greene and Cunningham, 2006) for AGNews. To evaluate generalizability, we use three sentiment classification datasets–IMDB, YELP, and Amazon Reviews (Zhang et al., 2015)–to purify victim SST-2 models through model merging. All datasets are downloaded from Hugging Face, which adheres to the Apache License 2.0.
Our method involves only one hyperparameter, $\alpha$, as described in  Equation˜1, which controls the preference between model utility and attack resistance during the model merging strategy search. We set $\alpha$ to 0.4 by default, which slightly favors seeking a more attack-resistant model. It can be adjusted to smaller values (e.g., 0.1) when a lower ASR is the primary target, and to larger values (e.g., 1.0) when high utility is demanded.
A.2 Computational Resources
We conduct experiments using three seeds on a single A100 GPU, and report the average scores. Running our method takes only 4 minutes on a single GPU for a 24-layer RoBERTa-large architecture.
BadNet InsertSent LWS HiddenKiller
[…]
Benign 4.1 95.9 2.2 95.9 12.8 95.9 16.5 95.9
Victim 100.0 96.0 100.0 96.3 98.0 95.4 96.5 95.7
Proxy Model (IMDB) 7.4 89.1 4.3 89.1 10.8 89.1 13.7 89.1
ONION 56.8 92.9 99.9 93.3 85.7 91.9 92.9 92.9
Z-Def. 4.6 96.1 1.8 95.6 97.3 95.3 35.7 95.4
PURE 0 50.9 0 50.9 0 50.9 0 50.9
TIES 99.9 95.7 100.0 95.8 93.5 95.2 88.8 95.3
DARE w/ TIES 99.3 96.0 100.0 96.2 96.4 95.7 92.7 95.8
WAG 84.4 94.8 60.1 95.2 58.8 94.8 56.2 92.5
SST-2 Ours (GMS) 4.5 91.6 1.9 92.5 9.7 91.7 10.4 91.2
Benign 29.0 84.9 31.0 84.9 45.6 84.9 53.6 84.9
Victim 99.9 85.2 100.0 85.0 94.5 85.1 100.0 85.0
Proxy Model (Twitter) 37.8 84.4 40.0 84.4 55.4 84.4 67.1 84.4
ONION 75.0 84.8 99.4 84.8 86.1 84.1 99.6 84.7
Z-Def. 29.7 85.6 30.3 85.3 93.5 85.3 53.9 85.7
PURE 82.2 75.2 87.5 75.2 76.1 79.1 100.0 72.1
TIES 36.2 84.8 38.2 84.8 57.5 84.5 65.3 84.5
DARE w/ TIES 95.6 86.3 77.8 86.0 90.5 85.9 86.8 86.3
WAG 52.0 85.0 48.6 84.6 63.1 84.5 68.2 85.0
OLID Ours (GMS) 32.1 85.0 28.0 84.2 52.2 84.4 64.2 84.9
Benign 12.3 87.6 12.6 87.6 26.4 87.6 36.9 87.6
Victim 100.0 89.4 100.0 90.3 96.0 89.0 99.9 89.4
Proxy Model (SNLI) 12.2 84.1 9.2 84.1 25.3 84.1 31.7 84.1
ONION 64.3 86.1 98.6 86.9 89.0 85.5 98.8 86.6
Z-Def. 11.1 88.3 11.6 89.7 92.2 89.1 50.6 89.7
PURE 33.3 33.8 33.3 33.8 33.3 33.8 33.3 33.8
TIES 91.6 89.2 94.5 89.5 80.7 89.2 88.6 90.0
DARE w/ TIES 93.5 90.5 100.0 91.4 93.9 89.9 99.4 90.5
WAG 71.0 88.7 60.3 88.5 77.9 88.0 80.5 88.8
MNLI Ours (GMS) 10.8 86.5 10.7 86.3 14.0 86.5 31.7 86.3
Benign 1.9 95.4 0.5 95.4 0.5 95.4 1.1 95.4
Victim 99.9 95.1 99.6 95.3 99.6 94.5 100.0 95.1
Proxy Model (BBCNews) 1.5 70.2 1.7 70.2 1.8 70.2 3.4 70.2
ONION 59.4 94.8 97.8 95.1 84.8 94.5 99.6 94.7
Z-Def. 1.6 95.3 0.4 95.3 97.9 96.1 100.0 95.0
PURE 2.8 86.3 3.0 85.4 6.5 85.0 9.4 84.6
TIES 99.9 94.6 99.6 94.4 97.7 95.8 100.0 94.4
DARE w/ TIES 99.9 95.2 99.6 95.4 97.8 96.5 100.0 95.2
WAG 92.7 94.1 97.8 94.4 78.0 93.9 90.9 94.3
AGNews Ours (GMS) 2.5 91.0 2.4 92.6 3.2 91.7 6.5 90.4
Table 8: Performance of our method compared to baselines under various backdoor attacks on the RoBERTa-large model, with each value averaged over three seeds.
A.3 Implementation details for baselines
We follow the open-source implementations for each baseline, and basically using their default hyper-parameters, while maintaining using the identical datasets, backdoor settings and the trained models for a fair comparison.
For the baseline DARE Yu et al. (2024), it first applies random parameter dropping and rescaling to the involved models with a specified drop rate, and then incorporates model merging techniques to combine the processed models. Various model merging methods can be integrated with DARE, and we choose TIES merging as a representative, as it demonstrates decent performance for encoder-based models (e.g., bert-base, roberta-base). Their method involves three tunable hyperparameters for encoder-based LMs, as outlined in Table 5 of their paper: drop rate, scaling term, and ratio to retain. We retained the original search space for drop rate and scaling term but expanded the ratio to retain from [0.1, 0.2, 0.3] to [0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8] for the roberta-large model to ensure more complete search.
For the baseline PURE (Zhao et al., 2024b), it first trains the victim model on a proxy clean dataset (from the same domain or a similar task, e.g., poisoned on IMDB and fine-tuned on SST-2) and prunes the attention heads that cause attention drifting between poisoned and clean texts. The pruned model is then fine-tuned further on the proxy clean dataset to normalize the remaining attention heads for purification. To ensure a fair comparison in the main experiments with the roberta-large model, we use the same poisoned and proxy datasets, e.g., poisoned on the SST-2 dataset with IMDB as the proxy clean dataset. We set the hyperparameter accuracy threshold (used to stop head pruning) to 0.90 for SST-2, MNLI, and AGNews, and 0.85 for OLID to prevent overly aggressive pruning. For the bert-base model, we follow the original implementation, including the use of the SST-2 dataset for fine-tuning SST-2 victim models and maintaining the default accuracy threshold of 0.85. While PURE uses the label flip rate (LFR) as its evaluation metric for backdoor defense (implanting triggers into test data while keeping the labels unchanged), we adopt the attack success rate (ASR) on label-flipped test data as our evaluation metric for a fair comparison with our method.
We adhere to the default implementations and hyperparameter settings for all other baseline methods.
Appendix B Additional results
BadNet InsertSent LWS HiddenKiller
[…]
Benign 8.5 92.9 3.5 92.9 22.2 92.9 17.3 92.9
Victim 100.0 92.9 100.0 92.2 98.1 91.5 95.9 91.8
Proxy Model (IMDB) 10.4 85.4 6.4 85.4 17.9 85.4 14.7 85.4
ONION 58.0 89.9 99.7 89.9 85.4 88.6 94.4 89.2
Z-Def. 8.3 92.5 1.8 91.9 97.4 90.9 38.7 91.5
PURE 30.1 92.0 10.7 91.5 66.4 91.1 28.7 91.3
TIES 99.2 92.7 98.2 92.4 91.2 92.3 86.8 92.6
DARE w/ TIES 100.0 93.0 90.8 88.9 92.6 92.7 94.1 90.6
WAG 74.3 91.9 70.7 91.7 73.5 91.7 60.3 91.6
SST-2 Ours (GMS) 10.8 86.8 9.8 88.8 27.4 88.0 19.5 86.8
Table 9: Performance of our method compared to baselines under various backdoor attacks on the BERT-base model, with each value averaged over three seeds. The victim model is trained on SST-2 dataset.
B.1 Performance across different model architectures and different dataset
Performance across different dataset. Due to space constraints, we present the complete performance results of our method compared to all baselines under various backdoor attacks on the RoBERTa-large model across all four datasets in Table˜2. The settings are consistent with those in Section˜5.2, with each score averaged over three runs using different seeds. For the OLID dataset, we specifically set $\alpha=0.1$ to enable more aggressive defense, as shown in Figure˜3. The top two ASR performances are highlighted. Across all datasets and backdoor attacks, our method consistently ranks among the top two in backdoor removal performance, with minimal harm to clean accuracy. Notably, under the two particularly challenging backdoor tasks, LWS and HiddenKiller, GMS achieves significant improvements (at least 25%) over all baselines. For instance, on SST-2, GMS reduces the ASR for LWS to 9.7%, compared to 58.8% for the next-best baseline, Z-Def.
While Z-Def achieves competitive performance with GMS in defending against BadNet and InsertSent, Z-Def is much less effective against LWS on both datasets. This is because Z-Def relies on lexical and syntactic features to detect outliers in the poisoned dataset, whereas LWS attacks subtly replace words with synonyms, effectively bypassing outlier detection. As for PURE, another competitive recent approach, we found its head-pruning step is highly unstable across both tasks and architectures. For example, it performs well on the BERT-base model in Table˜9 (following the settings reported in PURE (Zhao et al., 2024b)), but performs poorly on most tasks with RoBERTa-large. This instability seems to arise from PURE’s reliance on accuracy from the clean proxy dataset as the pruning stopping criterion. For well-trained models with high accuracy, a substantial number of heads are pruned, leading to a broken purified model (e.g., 0% ASR but random-guess-level CACC).
Additionally, although merging multiple models has been shown to be an effective backdoor defense, we observed that applying all merging baselines using a single proxy model is sub-optimal, aligning with the results in Arora et al. (2024). We found these methods prioritize preserving accuracy over removing backdoors. This behavior likely arises from the design of merging mechanisms, which are intended to preserve the performance of each merged model on downstream tasks. Consequently, backdoor tasks are treated equivalently to downstream tasks, exposing a persistent vulnerability.
Performance across different model architectures. We also evaluated our method on different architectures. Table˜9 shows the results on the BERT-base model. The poison rate remains at 20%, and each model is trained for three epochs. Scores are reported under $\alpha=0.4$ and averaged over three seeds: 1000, 2000, and 3000. Across all attacks, our method performs consistently well, particularly excelling against more challenging attack strategies. Notably, it achieves over a 39% advantage in defending against LWS and a 9% improvement over the second-best baseline for HiddenKiller.
B.2 Comparisons to other baseline defenses
Model purification baselines Fine-mixing Zhang et al. (2022) and Fine-purification Zhang et al. (2023) are two additional model purification baselines. We compare our method against their reported performance in their papers [^4] The code is not publicly available yet., strictly following their experimental settings. Specifically, we evaluate under the same conditions using two types of attacks (BadNet and InsertSent) and two model architectures (BERT-base and RoBERTa-base) on the AGNews dataset. As shown in Table˜10, our method consistently achieves better defense performance across all settings. Notably, against the more advanced InsertSent attack, our approach reduces the attack success rate (ASR) by at least 17% in absolute values compared to Fine-mixing and Fine-purification.
Architecture Methods BadNet InsertSent
CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$
BERT-base Fine-mixing 90.17 12.32 90.40 32.37
Fine-purifying 90.86 3.3 91.13 23.69
Ours 91.8 1.8 86.49 7.00
RoBERTa-base Fine-mixing 86.39 18.12 86.11 35.97
Fine-purifying 86.64 17.56 86.85 19.20
Ours 91.2 1.63 89.62 1.42
Table 10: Performance comparison of our method against Fine-mixing Zhang et al. (2022) and Fine-purifying Zhang et al. (2023) on AGNews dataset. The best results are in bold. Results in rows marked with are taken directly from the respective papers.
Poison sample detection with retraining: SEEP He et al. (2024)
While this paper focuses on the important but often overlooked post-deployment setting, it is natural to ask: if we had access to a highly effective data detection method capable of identifying most poisoned samples, how would our method compare? To explore this, we evaluate against SEEP He et al. (2024), a state-of-the-art poison detection method, on defending against BadNet attacks on the SST-2 dataset. As shown in Table˜11, although SEEP performs well under this relatively simple attack, our method–despite relying only on a naive heuristic for proxy data selection–achieves even better performance. As further discussed in Table˜6, incorporating more advanced prior knowledge (e.g., by using stronger detection methods to construct sub-proxy datasets) can further enhance the effectiveness of our approach.
[…]
RoBERTa-large SEEP with retraining 95.5 24.3
Ours 91.7 9.7
BERT-base SEEP with retraining 92.4 29.4
Ours 88.0 27.4
Table 11: Comparison of our method with retraining-based defenses on BERT-base and RoBERTa-large for SST-2 dataset.
B.3 Comparisons of all defenses under a more recent attack: BITE Yan et al. (2023)
BITE Yan et al. (2023) is a recent insertion-based textual backdoor attack that leverages label-biased tokens as stealthy triggers. Table˜12 shows the performance of our method and four baselines in defending against BITE on the SST-2 dataset. Our method consistently ranks among the top two in defense effectiveness, achieving comparable or even lower ASR than the benign model, while maintaining clean accuracy.
We exclude BITE from the main results in Section˜5.2 due to a peculiar issue: BITE tends to yield an unusually high ASR even on benign models. This undermines the reliability of ASR as an evaluation metric in this setting and makes it difficult to confidently interpret defense performance. We attribute this to BITE’s trigger selection process, which biases trigger tokens toward naturally label-correlated words, effectively exploiting model priors without requiring explicit poisoning.
Model Defense Method BERT-base RoBERTa-Large
[…]
Victim (BITE) Benign Model 92.9 41.4 95.9 38.3
No Defense 92.6 80.7 95.6 81.0
Defense Z-Defense 92.3 51.9 95.2 44.9
ONION 89.7 70.5 92.9 68.2
ABL 92.0 82.1 51.6 1.0
WAG 92.9 47.5 94.4 45.2
Ours 88.9 49.9 90.3 27.9
Table 12: Defense performance against the BITE Yan et al. (2023) attack on BERT-base and RoBERTa-large models for SST-2 dataset.
B.4 Performance under different poison rate: 20%, 10% and 5%
While the experiments in the main paper are conducted with a 20% poison rate to ensure fair comparison with the baselines, following the settings in (Dai et al., 2019; Qi et al., 2021b; Arora et al., 2024), i.e., 20% of the training samples are poisoned, we also evaluate lower poison rates of 10% and 5%, as shown in Table˜13. Our method is effective in model purification across all poison rates.
Victim Proxy Poison Rate Metric Result
[…]
Table 13: Results for SST-2 with varying poison rates.
Figure 2: Results of using different weights (alpha) on OLID dataset.
Figure 3: Results of using different weights (alpha) on AGNEWS dataset.
B.5 Stability: How sensitive is our method to the selection of hyperparameters?
We next show that GMS exhibits stable performance across a wide range of hyperparameters, in contrast to pruning-based and unlearning-based methods, which have been shown to be more sensitive to hyperparameter choices Liu et al. (2018); Wu et al. (2022) (as also observed in Table˜2). Our method involves two hyperparameters: $\alpha$ in Equation˜1 and stopping patience $T$, with default values of $\alpha=0.4$ and $T=5$ used in main results. Below, we justify these choices. As shown in Figure˜3 and Figure˜3, for most dataset and backdoor attack combinations, the default value $\alpha=0.4$ strikes a balanced trade-off between utility and backdoor defense strength. An exception is the HiddenKiller attack on the OLID dataset, where a more aggressive $\alpha=0.1$ is recommended. This is because OLID is a dataset collected from tweets, while HiddenKiller paraphrases the data using formal syntactic templates, drastically altering OLID’s language style. This change significantly impacts CACC, necessitating a higher weight for backdoor removal to ensure effective purification. For stopping patience, as shown in Figure˜4, the score computed in Equation˜1 generally follows a (weakly) monotonically decreasing trend. While a larger $T$ improves performance, we choose $T=5$ as a reasonable balance between performance and efficiency.
Figure 4: The search iteration history for four kinds of backdoors on SST-2 dataset.
Dataset Model CACC $\uparrow$
SST-2 Benign Victim Model 95.9
Proxy Model (IMDB) 88.7
Purified Model (GMS) 95.8
AGNews Benign Victim Model 95.2
Proxy Model (BBCNews) 78.1
Purified Model (GMS) 95.3
Table 14: CACC of benign victim models preserved.
[…]
Pure clean proxy dataset 88.3 4.7
Mixed poisoned proxy dataset 91.6 4.5
Table 15: Performance of our method with a clean vs. mixed proxy dataset, for the roberta-large model trained on the SST-2 dataset.
Methods
Metrics Stage WAG PURE Clean Ours
Before 94.72 93.12 95.76 91.86
[…]
Before 84.40 6.31 4.05 3.83
[…]
Table 16: Evaluating the robustness of different parameter purification methods towards the Retuning Attack (RA) using 0.1% of the original poisoned training set (which contains 20% poisoned data) for fine-tuning. We report the performance of each methods before and after the RA.
Figure 5: The optimal substitution strategy for defending against each backdoor attack for the roberta-large model trained on the SST-2 dataset. The green squares indicate the substituted modules.
B.6 If the victim model is benign, will our method affect its utility?
While all our previous experiments focus on malicious victim models, it is also important to consider scenarios where the victim model is benign. Does our method compromise the utility of a clean victim model? As shown in Table˜14, our results effectively dispel this concern. When applying GMS to a benign model, the CACC remains unaffected.
B.7 Can our method be effective with a fully clean proxy dataset?
While we primarily consider the practical scenario of receiving a mixed dataset, is our method still effective if the received dataset is entirely clean? As shown in Table˜15, when provided only with a clean proxy dataset—without any information about the poisoned dataset—GMS can still effectively remove backdoors. This is because our substitution strategy starts by fully removing backdoors and then gradually recovering task utility.
B.8 GMS remains effective as long as the proxy model is not backdoored with exactly the same attack and trigger.
Model Dataset Backdoored Method Before Substitution After Substitution
[…]
Victim SST-2 BadNet 95.6 100.0 95.3 4.5
Proxy SST-2 HiddenKiller 95.4 7.9
Victim SST-2 HiddenKiller 95.4 96.7 95.9 14.0
Proxy SST-2 BadNet 95.6 17.1
Victim SST-2 BadNet (Different triggers) 95.6 100.0 87.8 6.8
Proxy IMDB 87.7 8.6
Victim SST-2 BadNet (Same triggers) 95.6 100.0 95.8 100.0
Proxy IMDB 90.1 100.0
Table 17: Proxy models with implanted backdoors can still effectively purify victim models if the backdoor does not exactly match in both attack strategy and trigger.
An important question to consider is: What happens if the proxy model itself contains a backdoor? To investigate this, we considered scenarios where the victim and proxy models were either trained on the same dataset but with different backdoor attacks or on different datasets with the same attack.
From Table˜17, our findings indicate that GMS remains effective in backdoor removal as long as the proxy model’s backdoor is not an exact match to that of the victim model, i.e., both the same attack method and the same trigger.
B.9 Purification effectiveness as the poison ratio increases.
Due to space constraints, we included an incomplete version of Table˜7 in the main paper. Here, the full ratio table Table˜18 provides a clearer trend: (1) As the proportion of poisoned data in the proxy “clean” dataset increases, the CACC of the purified model slightly decreases. This suggests that under a default weight of $\alpha=0.4$ for $\mathcal{D}_{clean}$, a higher level of poisoning may introduce confusion, leading GMS to mistakenly remove some clean task-critical modules due to their reduced relative weight (see Appendix˜C for the analytical explanation). (2) However, even with poisoned ratios as high as 90% in SST-2 and 70% in AGNews, GMS remains effective in eliminating backdoor-related components and purifying the model. This effectiveness is mathematically justified in Appendix˜C.
SST-2 AGNews
Ratio $\rho$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$
0.00 93.58 3.38 91.76 2.42
0.10 93.58 3.38 91.76 2.42
0.20 91.06 4.28 91.59 3.61
0.30 91.51 3.15 91.59 3.61
0.40 89.11 5.18 91.59 3.61
0.50 89.11 5.18 91.59 3.61
0.60 89.11 5.18 88.67 1.42
0.70 89.11 5.18 88.67 1.42
0.80 89.11 5.18 94.59 99.75
0.90 89.11 5.18 94.83 99.82
Table 18: Purification performance as the ratio $\rho$ of poisoned data in the proxy “clean” dataset $\mathcal{D}_{clean}$ increases.
B.10 Robustness to Retuning Attacks
Recent works Min et al. (2024); Qi et al. (2024) have highlighted a critical pitfall in backdoor defenses: purified models with low ASR often fail to completely eliminate inserted backdoor features. A straightforward Retuning Attack (RA)–which involves fine-tuning a purified model on a small number of backdoored samples (e.g., 0.2% of the original poisoned data) mixed with benign samples (to maintain clean Accuracy on the main task)–can easily recover the backdoor in the purified model, as demonstrated in Table˜16.
Several related studies Qi et al. (2024); Chen et al. (2024b); Tarun et al. (2024) have sought to uncover the underlying reasons for this failure. A key observation is that the purified model’s parameters do not sufficiently deviate from those of the backdoored model. Specifically, the weight differences between a purified model and a (retrained) benign model are significantly larger than those between the purified and backdoored models. This suggests that purified and backdoored models are connected via a backdoor-related path in the loss landscape Min et al. (2024). Such vulnerabilities pose severe risks, as purified models are often deployed in various downstream applications, and even with backdoor defenses applied, attackers can easily re-trigger the backdoor in downstream tasks Min et al. (2024).
This challenge motivates us to propose a purification method aimed at breaking this backdoor-connected path. Intuitively, achieving this requires completely replacing "suspected" parameters inherited from the backdoored model, rather than pruning (Zhao et al., 2024b) or merging them with other parameters Arora et al. (2024); Yadav et al. (2023). Moreover, because parameters often interact within functional modules, modifying at the module level rather than the individual parameter level tends to remove backdoor features more effectively. As shown in Table˜16, when using 0.1% of the original poisoned training set to launch the Retuning Attack (RA) Min et al. (2024), the parameter-merging baseline, WAG Arora et al. (2024), is easily re-triggered. In contrast, our module-substitution method demonstrates greater robustness while maintaining utility, performing comparably to the clean model baseline (i.e., finetuning a clean model on the retuning dataset). We provide our baseline PURE with additional advantages by granting access to the full clean dataset of the victim task for finetuning, as obtaining a usable purified model otherwise proves challenging. While the recent method PURE also demonstrates competitive performance, it is worth noting that significant effort was required to find a configuration where its pruning step did not break the purified model. Even with these adjustments, our method still substantially outperforms PURE, achieving an ASR of 11.93% compared to 19.37%–nearly twice the robustness.
B.11 Examples of Searched Strategy
We applied our GMS defense to purify victim models compromised by various backdoor attacks across four datasets. Examples of the searched strategies from the SST-2 dataset are shown in Figure 5. While the obtained strategies vary, we speculate that they are transferable, as adapting the minimum replacement strategy identified for BadNet to the other three attacks yields effective defense results, as discussed in Section˜B.12.
BadNet InsertSent LWS HiddenKiller
[…]
Benign 4.1 95.9 2.2 95.9 12.8 95.9 16.5 95.9
Victim 100.0 96.0 100.0 96.3 98.0 95.4 96.5 95.7
Proxy Model (IMDB) 7.4 89.1 4.3 89.1 10.8 89.1 13.7 89.1
GMS 4.5 91.6 1.9 92.5 9.7 91.7 10.4 91.2
SST-2 SST-2 BadNet Strategy 4.5 91.6 9.8 93.0 22.3 92.8 13.2 90.3
Benign 29.0 84.9 31.0 84.9 45.6 84.9 53.6 84.9
Victim 99.9 85.2 100.0 85.0 94.5 85.1 100.0 85.0
Proxy Model (Twitter) 37.8 84.4 40.0 84.4 55.4 84.4 67.1 84.4
GMS 32.1 85.0 28.0 84.2 52.2 84.4 64.2 84.9
OLID OLID BadNet Strategy 33.6 85.0 37.4 84.6 56.5 84.3 65.0 84.7
Benign 12.3 87.6 12.6 87.6 26.4 87.6 36.9 87.6
Victim 100.0 89.4 100.0 90.3 96.0 89.0 99.9 89.4
Proxy Model (SNLI) 12.2 84.1 9.2 84.1 25.3 84.1 31.7 84.1
GMS 10.8 86.5 10.7 86.3 14.0 86.5 31.7 86.3
MNLI MNLI BadNet Strategy 12.2 86.3 10.8 86.3 27.0 85.2 35.2 86.4
Benign 1.9 95.4 0.5 95.4 0.5 95.4 1.1 95.4
Victim 99.9 95.1 99.6 95.3 99.6 94.5 100.0 95.1
Proxy Model (BBCNews) 1.5 70.2 1.7 70.2 1.8 70.2 3.4 70.2
GMS 2.5 91.0 2.4 92.6 3.2 91.7 6.5 90.4
AGNews AGNews BadNet Strategy 1.4 90.7 1.2 89.8 2.6 90.1 12.0 90.2
Table 19: The performance of transferring strategy searched based on BadNet to other attacks.
B.12 Transferability of substitution strategies
Interestingly, we find that GMS exhibits strong transferability: a substitution strategy optimized for one attack often generalizes to others, enabling defense even without access to data. Visually, the module substitution strategies ($S_{best}$) identified by GMS across different backdoor attacks (examples in Figure˜5 in Section˜B.11), though distinct, share many similar patterns across attacks.
In Table˜19, for each task dataset, we apply the substitution strategy found on BadNet (replacing the fewest modules and layers as shown in Figure˜5) to to determine which modules to substitute in the three other victim models under different attacks, replacing them with modules from their corresponding proxy models, i.e., directly executing Line 24 in Algorithm˜1 using $S_{badnet}$. Compared with the complete GMS (i.e., searching for the optimal strategies $S_{best}$ for each attack), the transferred strategy consistently performs comparably. This demonstrates that for defenders who only have access to the victim model and are unaware of the victim dataset, a universal GMS defense strategy exists for each task that can effectively defend against multiple attacks.
Appendix C Constraints on the proxy datasets $\mathcal{D}_{clean}$ and $\mathcal{D}_{poison}$
Table˜7 and Table˜6 illustrates that our method is quite robust to the construction of the proxy datasets $\mathcal{D}_{clean}$ and $\mathcal{D}_{poison}$, e.g., even random sampling (for $\mathcal{D}_{clean}$) and inaccurate heuristics (for $\mathcal{D}_{poison}$) can yield effective defense. Natural questions arise: Are there any constraints on the proxy datasets? Could the random sampling strategy for constructing $\mathcal{D}_{clean}$ fail if $\mathcal{D}_{poison}$ is highly inaccurate? We next establish the constraints and demonstrate that, in most cases, even when only half of the extracted samples in $\mathcal{D}_{poison}$ are poisoned, our approach can still use random sampling for $\mathcal{D}_{clean}$ to effectively purify the model.
Let the constant values $s_{asr}(M_{v})$ and $s_{acc}(M_{v})$ in Equation˜2 and Equation˜3 be denoted as $c_{1}$ and $c_{2}$, respectively, for simplicity. We use ASR and ACC to represent the ground-truth attack success rate and clean accuracy of a model (that can be measured on oracle poisoned and clean test sets), respectively.
Assuming we have two imperfect proxy datasets: a proxy “clean” dataset that contains a proportion $\rho$ of poisoned data and a proxy “poisoned” dataset contains a proportion $1-\lambda$ of the non-poisoned data. Denote the poisoned and clean data distributions as $\mathcal{P}$ and $\mathcal{C}$, respectively.
Then, by the rule of total probability and definition in Section˜3, we obtain:
$s_{acc}(M)$ $=\mathbb{E}_{x\sim\mathcal{D}_{clean}}[\mathbf{1}(M(x)=y)]$
$=p(x\in\mathcal{P}\mid x\sim\mathcal{D}_{clean})$
$\quad\cdot p(M(x)=y\mid x\in\mathcal{P},x\sim\mathcal{D}_{clean})$
$\quad+p(x\in\mathcal{C}\mid x\sim\mathcal{D}_{clean})$
$\quad\cdot p(M(x)=y\mid x\in\mathcal{C},x\sim\mathcal{D}_{clean})$
$=\rho\text{ASR}+(1-\rho)\text{ACC}.$
Similarly, for $x\sim\mathcal{D}_{poison}$, we have:
$s_{asr}(M)$ $=\lambda\cdot\text{ASR}+(1-\lambda)\cdot\text{ACC}.$
Substituting these into Equation˜1 gives us $(1-\alpha)c_{1}+\alpha(1-c_{2})+\left(\alpha\rho+\alpha\lambda-\lambda\right)\text{ASR}+\left(2\alpha-1-\alpha\rho-\alpha\lambda+\lambda\right)\text{ACC}$. From this, the constraints ensuring effective purification are:
$\left\{\begin{array}[]{ll}\alpha\rho<(1-\alpha)\lambda,\\ \alpha(1-\rho)>(1-\alpha)(1-\lambda).\end{array}\right$. (6)
That is, as long as the clean portion in $\mathcal{D}_{clean}$ receives more attention than that in $\mathcal{D}_{poison}$, our method will effectively remove backdoor modules while preserving utility; the same applies to the poisoned portion in reverse.
To further illustrate, for an accurate proxy dataset, the constraints are simplified to $\alpha(1+\rho)-1<0$ and $\alpha(1-\rho)>0$. Thus, if $\rho<1$, setting $\alpha$ such that $0<\alpha<\frac{1}{\rho+1}$ can ensure effective purification.
In our experiments, $\alpha$ is set to 0.4 by default. Then, the requirement becomes $\rho<\frac{3}{2}\lambda-\frac{1}{2}$. This implies that even with a random-guess backdoor data detection method, we can tolerate up to $25\%$ poisoned data in $\mathcal{D}_{clean}$. Given that real-world poisoning rates are usually low (e.g., below $20\%$ or even $1\%$), random sampling for the “clean” proxy dataset remains highly feasible in practice.
Appendix D AI Assistants
We use ChatGPT/Gemini for writing and formatting supports, including grammar checks, improving the clarity of figure and table captions, and other surface-level edits.